### v1

# Sparsity May Be All You Need: Sparse Random Parameter Adaptation

Abstract
Full fine-tuning of large language models for alignment and task adaptation has become prohibitively expensive as models have grown in size. Parameter-Efficient Fine-Tuning (PEFT) methods aim at significantly reducing the computational and memory resources needed for fine-tuning these models by only training on a small number of parameters instead of all model parameters. Currently, the most popular PEFT method is the Low-Rank Adaptation (LoRA), which freezes the parameters of the model to be fine-tuned and introduces a small set of trainable parameters in the form of low-rank matrices. We propose simply reducing the number of trainable parameters by randomly selecting a small proportion of the model parameters to train on. In this paper, we compare the efficiency and performance of our proposed approach with PEFT methods, including LoRA, as well as full parameter fine-tuning.
Sparsity May Be All You Need:
Sparse Random Parameter Adaptation
Jesus Rios,   Pierre Dognin,   Ronny Luss,   Karthikeyan Natesan Ramamurthy IBM Research {jriosal, pdognin, rluss, knatesa}@us.ibm.com
1 Introduction
It has become common practice to train application-ready language models in two phases (Radford et al., 2018; Kenton and Toutanova, 2019): first, the model is pre-trained on a very large and general corpus of (unlabeled) text; then further trained (or fine-tuned) on a smaller specific set of examples demonstrating the intended behavior for a particular application, such as instruction following (Ouyang et al., 2022b).
Overall, supervised fine-tuning (SFT) requires less computational resources than pre-training (PT) due to the significantly smaller size of the training set as well as the typical use of early stopping to deal with overfitting. This means orders-of-magnitude less gradient computations and parameter updates are needed during SFT compared to PT. However, a major drawback is that memory requirements remain the same, unless a parameter-efficient fine-tuning (PEFT) technique is used. The main memory bottleneck during training is the number of trainable parameters, since additional memory must be allocated for their gradients and other per-parameter statistics needed by the optimizer. The idea behind PEFT (Lialin et al., 2023) is to significantly reduce the number of trainable parameters during fine-tuning while maintaining performance.
Figure 1: The proposed SpaRTA method which randomly chooses a subset of parameter indices vector $\phi$ from pre-trained model parameters $\theta_{\text{PT}}$ and updates them via adapter $\Delta_{\phi}$.
Low Rank Adaptation (LoRA), first introduced by Hu et al. (2022), currently remains the most popular PEFT technique. LoRA freezes all the pre-trained model parameters $\theta_{\text{PT}}$ and introduces trainable low-rank matrices (e.g., $B$, $A$) for a pre-selected subset of matrix tensors to represent the changes ($\Delta=BA$) needed for adapting the model to a new task. The adapted model parameters are given by $\theta_{\text{PT}}+\Delta$. Memory and computational efficiency are achieved by optimizing only over the parameters of these newly added, but significantly smaller, matrices.
The success of LoRA begs one to ask what properties make this method perform well. Is the low-rank structure critical, i.e., does $\Delta$ need to be low-rank? Is it sufficient to constrain $\Delta$ to be low dimensional? A main goal of this paper is to investigate these questions. An abundance of research is going into new methods for structured $\Delta$ (see Section 3 below) and novel directions into unstructured methods for fine-tuning could open avenues in areas such as model merging Wortsman et al. (2022); Matena and Raffel (2022) or pluralistic alignment Feng et al. (2024).
In this work, we propose a different approach where $\Delta$ is not factorized into low rank matrices but rather chosen to be a random subset of the model parameters. This Sparse Random parameTer Adaptation (SpaRTA) method imposes a sparsity constraint on the adapter that can be easily controlled. By changing the desired sparsity, one can change the number of adaptation parameters. Regardless of how the selected parameters are sampled from the model parameters, subsequent updates only affect these parameters. This sparsity constraint and randomness of selected parameters is in contrast to techniques such as LoRA that effectively affect all parameters in $\theta_{\text{PT}}$. See Figure 1 for an illustration of the method. Generally, one samples $m$ parameters from the pre-trained model $\theta_{\text{PT}}$, stores the indices of these parameters in $\phi$, and uses adapter $\Delta_{\phi}$ to fine-tune $\theta_{\text{PT}}$.
To investigate the performance of SpaRTA, we build adapters with different sparsity levels and test on the GLUE classification tasks (Wang et al., 2019) as well as the IMDB dataset (Maas et al., 2011). SpaRTA is compared to a fully fine-tuned model (Full FT), and PEFT approaches including LoRA. SpaRTA is quite competitive compared to LoRA given that it only modifies a small sparse number of model parameters.
2 Motivation
Yu et al. (2024) look at the differences between a language model’s parameters before and after fine-tuning, and demonstrate empirically that it is possible to randomly drop (i.e., set to zero) up to 99% of these parameter changes without significantly affecting the model performance. The sparsity in the parameter changes represented by the tensor $\Delta$ is known as $\Delta$-sparsity.
This motivates our approach, SpaRTA, which produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters: SpaRTA randomly selects the (scalar) parameters to train and freezes the remaining parameters (i.e., setting the corresponding $\Delta$ values to zero). This parameter sparsity also helps in reducing overfitting, as pre-trained models typically have enough capacity to learn the often limited amount of (labeled) data used for fine-tuning. There is no guarantee of $\Delta$-sparsity in LoRA, but this is a desired property since it reduces parameter interference (Yadav et al., 2023) when merging fine-tuned, task-specific models into a single one that can perform all tasks simultaneously.
One byproduct of $\Delta$-sparsity is sparse $\Delta$ updates which reduce gradient computations during training, ensure fast merge back of the sparse $\Delta$ into the model for inference, and ultimately same inference cost as the original model (like LoRA). The properties of low memory, low computation needs in training, identical inference time as the original model are all quite desirable for an adaptation technique. SpaRTA has all these properties plus the unique added benefit of producing only sparse changes in a small number of the parameters of the original model that can be beneficial for merging multiple SpaRTA adapters.
2.1 Is low-rank necessary for $\Delta$?
In Appendix A, we show that the weight matrix changes (i.e., $\Delta$) during full parameter fine-tuning are, in fact, not generally low-rank for capable models such as gemma-2b and mistral-7b. This indicates that LoRA works, not because of the low-rank constraint particularly, but rather due to the capacity reduction achieved since fine-tuning with LoRA is done with limited training data. Such insight also hints that any constraint which reduces the capacity of the original model could perform competitively, motivating our approach that selects a small number of parameters of the original model to be updated during training.
3 Related work
The last few years has seen many advances in fine-tuning Large Language Models (LLMs). Perhaps the most well-known method (also most used in practice) is LoRA Hu et al. (2022), which has spurred many variants including DoRA – which adapts only the directions of pre-trained weights Liu et al. (2024), VeRA – which shares low-rank matrices across layers Kopiczko et al. (2024), AdaLoRA – which adaptively allocates parameter budgets among weight matrices according to their importance Zhang et al. (2023), and SoRA – which dynamically adjusts the intrinsic rank during the adaptation Ding et al. (2023). This set of methods each have a different structure to the $\Delta$ being optimized with the commonality being the training of some function of low-rank matrices. For example, further expanding on VeRA Kopiczko et al. (2024), the authors propose to learn $\Delta=DBEA$ where $A$ and $B$ are random low-rank matrices and $D$ and $E$ are diagonal matrices to be trained.
Beyond adding structured parameters is the concept of fine-tuning a small subset of the total parameters of the model, i.e., sparse fine-tuning, where one must first decide which subset of parameters to fine-tune (similar to deciding which parameters in each layer to add adapters to in LoRA). Ansell et al. (2024) start with an initial set of parameters and offer procedures to drop (according to parameter magnitude) and grow (according to gradient) the set, i.e., they learn the set of parameters to train. Alternatively, Ma et al. (2024) focus on the sparsity of neuron activations during training by pre-computing neuron importance scores and only including important neurons in computations during training. Ansell et al. (2022) first fine-tune on all parameters, select the parameters that change the most, and then fine-tune again from scratch on the selected parameters. In contrast to these, our proposed SpaRTA method produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters chosen completely at random.
Yet another related direction is that of compression which results in sparse models; algorithms in this genre take a dense fine-tuned model with the goal of compressing it while maintaining performance. Compression (see Zhu et al. (2024) for a survey) could be accomplished by quantization, low-rank approximation, pruning (i.e., removing neurons, attention heads, or even layers), or distillation. The focus of this paper, however, is on fine-tuning dense models rather than learning sparse models as in Liu et al. (2023).
4 SpaRTA: Adapting a Random Subset of Model Parameters
Suppose the vectorized parameters of the pre-trained language model are $\theta_{\text{PT}}\in\mathbb{R}^{n}$ where $n$ is the number of parameters, and full parameter fine-tuning (FT) of the model is performed with a labeled dataset characterizing a task. FT typically updates $\theta_{\text{PT}}$ using a stochastic first-order gradient-based optimization algorithm (e.g., Adam from Kingma and Ba (2015)) to maximize the conditional probability of the labels in the training dataset under the model. The FT model is then given by $\theta_{\text{FT}}=\theta_{\text{PT}}+\Delta_{\text{FT}}$ where $\Delta_{\text{FT}}\in\mathbb{R}^{n}$.
This has two drawbacks in terms of memory efficiency. First, the optimizer state can be large, e.g., the state of the Adam optimizer is $4$ times as large as the parameter space as it includes current parameter values, a moving average of gradients (per parameter), etc. Second, storing a new FT model takes as much memory as the PT model which could be an issue when many task-specific models are requested.
SpaRTA proposes to randomly select a small subset of the model parameters to optimize while freezing the rest. The model parameters $\theta\in\mathbb{R}^{n}$ are partitioned into trainable $\theta_{\text{T}}\in\mathbb{R}^{m}$ and frozen $\theta_{\text{F}}\in\mathbb{R}^{n-m}$ ones, with $m\ll n$ being the number of parameters selected. This allows our approach to have a drastically lower memory footprint than FT (similar to LoRA) by reducing the size of the optimizer state as well as faster training by reducing the number of gradients to compute.
In our SpaRTA’s implementation, we introduce: (i) non-trainable indices $\phi\in\mathbb{R}^{m}$ containing the indices of the randomly selected elements in $\theta_{\text{PT}}$, and (ii) trainable parameters $\Delta_{\phi}\in\mathbb{R}^{m}$ representing the subset of $\Delta$ that SpaRTA learns at indices $\phi$. The pseudocode for SpaRTA is given in Algorithm 1.
Input: Pre-trained model with $\theta_{\text{PT}}\in\mathbb{R}^{n}$
Input: Task labeled dataset $\mathcal{D}$
Sample $\phi$: Indices of parameters to be optimized
Initialize: $\Delta_{\phi}=0$
while validation loss has not converged do
      1. Merge: $\theta_{\phi}=\theta_{\phi}+\Delta_{\phi}$
      2. Compute loss: Continue forward pass from $\theta_{\phi}$ using a batch of labeled data from $\mathcal{D}$
      3. Compute gradients with respect to $\Delta_{\phi}$ using backpropagation
      4. Unmerge: $\theta_{\phi}=\theta_{\phi}-\Delta_{\phi}$, without recording this operation in the computational graph
      5. Update $\Delta_{\phi}$ with the Adam Optimizer
Output: $\phi,\,\Delta_{\phi}$
Algorithm 1 SpaRTA
Whereas optimizing LoRA requires computing gradients with respect to the $A$ and $B$’s used to compute $\Delta$ (which constitute additional parameters, independent from $\theta_{\text{PT}}$), SpaRTA requires computing gradients with respect to $\Delta_{\phi}$, the changes over a subset of parameters chosen from $\theta_{\text{PT}}$ and indexed by $\phi$. This nuance is why LoRA can trained with out-of-the-box optimizers, but SpaRTA requires a different implementation. For example, Step 4 in Algorithm 1 must take into account the computational graph so as to not interfere with the forward pass to be computed at the following iteration. Specifically, this unmerge operation must not be recorded in the computational graph.
Generating the index set $\phi$ is done as follows. For each $i\in\{1,\ldots,n\}$, SpaRTA samples $x\sim\mbox{Bernoulli}(m/n)$ and includes $i\in\phi$ if $x=1$. Hence, SpaRTA uses $m$ trainable parameters in expectation. Step 5 also depends on the specific optimizer and settings being used; SpaRTA uses Adam in our implementation. Finally, inference is performed, similarly to LoRA, by merging the fine-tuned $\Delta$ into $\theta_{\text{PT}}$ and then making the necessary forward passes. Thus, SpaRTA does not introduce any additional inference latency when deploying the resulting fine-tuned model.
5 Memory Usage
Recall that SpaRTA freezes $n-m$ ($m\ll n$) of the model parameters. We define sparsity as $s=1-m/n\in(0,1)$, the percentage of frozen model parameters (e.g., if $1\%$ of parameters are trainable, then the sparsity is $99\%$). Subsequently, density is defined as $k=m/n=1-s$, the percentage of trainable model parameters. In practice, for a chosen sparsity $s$, one can freeze a model parameter with probability $s$, expecting a total sparsity percentage of $s$ over all model parameters. Thus, in expectation, $k=m/n$ percent of the model parameters are chosen as trainable, for a total of $n\,k=m$ trainable parameters.
For SpaRTA, only $\Delta_{\phi}$ (of size $m$) is trainable, which is significantly smaller than the total number of trainable parameters for FT since $m\ll n$. However, the indices of these randomly chosen model parameters must be recorded into $\phi$, adding to the memory requirements. Indices can be stored in $16$-bit integers for all the Transformer models (Vaswani et al., 2017) considered in this paper.
SpaRTA sparsifies neither the model head (which is kept fully trainable) nor the embeddings (which is kept frozen during training). The parameters in Transformer networks consist of bias vectors and two-dimensional weight matrices. Storing indices for all these trainable parameters would require at most $m\,(2\times 16)$ bits of memory ($m$ parameters, two integers to index a 2-dimensional matrix, 16 bits per integer).
The values in $\Delta_{\phi}$ are of the same type as the model parameters, e.g., using $16$-bit brain floating-point (bfloat16) tensors. Thus, SpaRTA requires up to $m\,(2\times 16+16)=3m\times 16$ bits of extra memory to specify the index $\phi$ and delta $\Delta_{\phi}$ tensors. That is $3\,k$ times more memory than the original model, which requires just $n\times 16$ bits for storing its parameters. For instance, using SpaRTA on a model with sparsity 80%, 90%, 95% and 99% would require up to 60%, 30%, 15% and 3% more memory, respectively.
We next analyze [^1] This analysis does not include memory requirements associated with model buffers (e.g., those used to track running statistics for layer normalization) because they are relatively small and the same for both the Full FT and SpaRTA. SpaRTA’s memory savings during training. Optimizing the full set of model parameters (FT) using Adam (Kingma and Ba, 2015) requires memory for the parameters, their gradients with their (adaptive) first and second moments estimates. This requires $4\,n\times 16$ bits of memory when using a bfloat16 representation.
In contrast, SpaRTA only optimizes $\Delta_{\phi}$, requiring a total of $m\,(4\times 16+2\times 16)+n\times 16$ bits of memory: (i) $m\,(4\times 16)$ bits for the statistics needed by Adam to optimize $\Delta_{\phi}$ of size $m$, (ii) $m\,(2\times 16)$ bits for $\phi$ with the indices identifying the model parameters associated with $\Delta_{\phi}$, and (iii) the PT model parameters ($n\times 16$ bits).
Memory savings appear if and only if
$m\,(4\!\times\!16+2\!\times\!16)+n\times 16$ $<4n\!\times\!16,$ (1)
$kn\,(6\!\times\!16)$ $<3n\!\times\!16,$
$k$ $<\frac{1}{2},$
that is, SpaRTA is a real PEFT method iff $k<0.5$, i.e., a sparsity higher than 50% is required ($s\!>\!0.5$), making less than 50% of the model parameters trainable. For instance, using SpaRTA on a model with sparsity $s=$ 80%, 90%, 95%, and 99% requires 45%, 60%, 67.5% and 73.5% less memory than full parameter FT, respectively, see Table 1.
Model storage Full FT SpaRTA $n$ (on disk) 0% 50% 80% 90% 95% 99% 2B 4 16 16 8.8 6.4 5.2 4.2 7B 14 56 56 30.8 22.4 18.2 14.8
[…]
6 Experimental Setup
We next detail the framework for conducting the experiments in Section 7. Motivation is first given for the tasks followed by a description of models to be used and how these models are adapted (with manipulations) for the desired tasks.
6.1 Tasks
The main experimental focus is on Natural Language Understanding (NLU) tasks, specifically sequence classification, which involves classifying natural language sequences into a given number of classes. NLU tasks are easier to evaluate than Natural Language Generation (NLG) tasks, as one can use simple metrics (Accuracy, F1, Recall, Precision, etc.) that can avoid the inherent ambiguity of evaluating more general generative tasks. While SpaRTA is also applicable to NLG tasks, they are not used in the following demonstrations due to the challenges associated with their evaluation, typically requiring human judgments as the gold standard measure of performance. Human evaluations can be expensive and time consuming, do not scale easily, and are not as objective as NLU evaluations.
6.2 Language Models
Starting with available open-weight language models, the goal is to adapt them to perform sequence classification. Two types of trained models are used: base and instruction-tuned models, where the latter have additionally been trained to follow users’ intent when prompted, such as answering questions or responding to instructions. Specifically, we consider the following language models: gemma-2b and gemma-2b-it from the Gemma (Team et al., 2024) family, and mistral-7b and mistral-7b-it from the Mistral [^2] We use v0.3 for both models. (Jiang et al., 2023) family, developed by Google and Mistral AI respectively. The "it" suffix refers to instruction-tuned models. They are of particular interest for our experiments as they will show results for models with two different numbers of parameters. All models are readily available text-to-text, decoder-only transformer models with open weights, downloadable from Hugging Face [^3] https://huggingface.co/(HF). Note that when using an instruction-following (it) model, inputs are formatted according to the conventions established when training the model on instructions.
6.3 Task Adaptation
When using a base model for sequence classification, the raw sequences to be classified are input directly into the model. However, when using an instruction model, these sequences are first wrapped into a classification-specific instruction. For example, a possible instruction could be: “Determine if the following sentence has a positive sentiment. Respond Yes or No.”, followed by the sequence itself to be classified.
A (generative) pre-trained Transformer model with a decoder-only architecture has a head that transforms the final hidden state of each token in the input sequence into vocabulary logits. Efficiently adapting this model for sequence classification requires the swap of this head for a sequence classification head, which uses only the final hidden state of the last token in the input sequence $h\in\mathbb{R}^{d}$ to do the classification. This reduces the parameter size of the head, which is just a linear layer with no bias applied to a token’s final hidden state, from a weight matrix $W\in\mathbb{R}^{v\times d}$ to $W\in\mathbb{R}^{c\times d}$, where $d$ is the dimension of the model’s hidden states (e.g., 2,048 and 4,096 for the gemma-2b and mistral-7b models respectively), $v$ is the number of tokens in the vocabulary (e.g., 256,000 for gemma-2b or 32,768 for mistral-7b), and $c$ is the number of classes in the individual task (e.g. $2$ to $3$ in the experiments that follow). With this small change, the model outputs classification probabilities for each whole input sequence through
$p=\operatorname*{softmax}(h\,W^{T})\in\mathbb{R}^{c}.$ (2)
The classification head of our base pre-trained models (i.e., gemma-2b and mistral-7b) is initialized with random weights, sampled from a normal distribution $\mathcal{N}(0,\sigma^{2})$, with $\sigma$ chosen such that the model at initialization predicts either class with equal probability. This is achieved with $\sigma=0.06/\sqrt{d}$. With this, the loss starts at approximately $-\log(1/2)\approx 0.693$ and the accuracy at $\approx 50\%$ for a balanced evaluation dataset.
Regarding instruction-tuned models (i.e., gemma-2b-it and mistral-7b-it), the original vocabulary weights are rather reused when initializing their classification heads. To do so requires to first identify the tokens in the vocabulary that the model is expected to use for classification following the instruction. For example, these could be the tokens associated with a “Yes” or “No” response by the model. The embeddings in the original model (head) associated with those classification tokens are extracted and used to initialize the classification head. While many models tie their vocabulary heads to their tokens embedding matrices, these new classification heads are never tied to the model’s input embedding matrix.
7 Experimental Results
Efficacy of SpaRTA is demonstrated empirically by comparing it to the following baselines:
  • Full parameter Fine-Tuning (Full FT): all model parameters are optimized.
  • LoRA: Only the additional matrices representing low-rank adapters are optimized.
  • Head: Only the classification head is optimized; all other parameters remain frozen.
This enables the exploration of the performance of SpaRTA on a range of sparsity levels, varying the number of trainable parameters. A complete set of results can be found in Appendix D accompanied by a detailed description of the training setup in Appendices C and E.
7.1 Datasets
Each sequence classification task considered in our experiments is given by a dataset of labeled examples. Table 2 summarizes these datasets and their splits for training, development, and testing for both the GLUE classification tasks (Wang et al., 2019) and IMDB (Maas et al., 2011). See Appendix B for more detailed descriptions.
Dataset Classes Train Dev Test
COLA 2 7,551 1,000 1,043
MNLI 3 100,000 10,000 19,647
MRPC 2 3,668 408 1725
QNLI 2 99,725 5,000 5,463
QQP 2 100,000 5,000 40,430
RTE 2 2,182 300 277
SST-2 2 66,349 1,000 872
IMDB 2 25,000 5,000 20,000
Table 2: Sequence classification datasets. Training sets limited to 100K samples. Training samples with > 256 tokens are removed (here using gemma-2b tokenizer, with mistral-7b tokenizer in Appendix B).
7.2 IMDB
gemma-2b gemma-2b-it mistral-7b mistral-7b-it Adaptation density (%) loss acc. loss acc. loss acc. loss acc. Full FT 100% 0.092 96.9% 0.107 96.3% 0.080 97.4% 0.080 97.3% SpaRTA 5% 0.096 96.7% 0.105 96.3% 0.084 97.2% 0.081 97.5% SpaRTA 0.5% 0.096 96.7% 0.103 96.4% 0.087 96.9% 0.087 97.0% SpaRTA 0.05% 0.106 96.3% 0.114 96.2% 0.080 97.3% 0.076 97.4% LoRA $\approx$0.050.101 96.5% 0.113 96.2% 0.086 97.1% 0.081 97.1% Head $\approx$2e-40.165 93.6% 0.194 92.7% 0.139 94.9% 0.117 95.8%
Table 3: Test loss and accuracy of models adapted to the IMDB dataset with different fine-tuning methods. For training details see Table 12.
Table 3 shows the results for the IMDB classification task where each model is asked to rate a review as positive or negative. Each model is fine-tuned using adaptation techniques including: (i) Full FT where all model parameters are updated (100% density); (ii) SpaRTA for different orders of magnitude of density levels 5%, 0.5%, $\approx$0.05– the last one allowing for SpaRTA to have close to the same number of trainable parameters as for LoRA; (iii) LoRA with rank $r=8$ equivalent to $\approx 0.05\%$ of trainable parameters compared to the model full parameter size; (iv) Head adaptation where only the classification head is updated ($\approx$2e-4density).
Model: gemma-2-2b-it QNLI RTE SST2 QQP MNLI MRPC COLA loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc Full FT (100%) 0.46 81.2 0.33 87.4 0.18 93.3 0.36 85.5 0.26 88.9 0.12 96.1 0.36 63.6 SpaRTA (5%) 0.45 79.8 0.33 87.6 0.18 93.0 0.40 81.3 0.24 89.9 0.21 95.6 0.41 57.2 SpaRTA (0.5%) 0.55 72.9 0.37 86.1 0.25 89.8 0.34 85.4 0.24 89.9 0.13 95.6 0.39 56.9 SpaRTA (0.037%) 0.45 79.8 0.38 85.5 0.18 93.0 0.40 82.3 0.27 88.6 0.12 95.8 0.45 56.5 LoRA (0.037%) 0.46 78.3 0.33 87.5 0.24 90.4 0.36 84.8 0.26 89.0 0.13 95.5 0.39 59.9 Head (1.63e-4%) 2.60 47.3 0.73 67.4 0.96 37.1 0.55 74.2 0.66 64.9 0.55 84.2 0.62 18.0
Table 4: Test loss and accuracy of model google/gemma-2-2b-it adapted to glue datasets with different fine-tuning methods. For training details see Table 13.
In Table 3, the results for adaptation techniques (rows) are sorted in order of descending density as to show the impact of decreasing the number of trainable parameters on the overall performance. For gemma-2b and gemma-2b-it, Full FT adaptation gives strong test loss and accuracy numbers, often the best available – this is expected since all model parameters are fine-tuned. SpaRTA results at 5% density are close to Full FT, even matching them for gemma-2b-it. As the density decreases by orders of magnitude, the results slowly degrade. For $\approx$0.05%, gemma models’ performances are slightly behind or matching performance from LoRA. Head adaptation provides the worse results as one would expect since it has so few parameters to work with.
For mistral models, the trend is similar with Full FT adaptation showing strong loss and accuracy numbers. SpaRTA performs well, matching and even improving upon Full FT with a density of 0.05%. SpaRTA even improves over LoRA for both mistral-7b and mistral-7b-it at this low density. Once again, Head adaptation gives the worst results. Overall these results are encouraging and show that SpaRTA is competitive and provides similar performance to LoRA for low densities.
7.3 GLUE
Model: Mistral-7B-Instruct-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc Full FT (100%) 0.32 91.7 0.26 91.3 0.14 95.7 0.33 88.6 0.23 90.3 0.11 96.7 0.32 70.5 SpaRTA (5%) 0.30 89.9 0.23 91.4 0.12 95.8 0.28 88.9 0.23 90.5 0.10 96.7 0.30 72.1 SpaRTA (0.5%) 0.29 89.5 0.23 91.6 0.13 95.2 0.29 88.6 0.25 89.2 0.10 96.6 0.32 70.3 SpaRTA (0.048%) 0.27 89.9 0.24 91.1 0.12 95.9 0.31 87.6 0.25 89.3 0.10 96.6 0.32 69.9 LoRA (0.048%) 0.32 88.8 0.26 90.4 0.12 95.6 0.32 87.2 0.27 88.5 0.11 96.9 0.30 71.3 Head (1.15e-4%) 0.31 86.3 0.35 87.0 0.23 90.4 0.35 84.5 0.32 85.5 0.15 93.5 0.36 62.1
Table 5: Test loss and accuracy of the mistralai/Mistral-7B-Instruct-v0.3 model adapted to glue datasets with different fine-tuning methods. For training details see Table 13.
Table 4 compares the results of gemma-2b-it on 7 of the GLUE tasks for the same adaptation methods as in Table 3. An equivalent set of results for the gemma-2b base model is given in Table 10, see Appendix D. Here again, adaptation results (rows) are ordered in descending order of density with SpaRTA evaluated for different orders of magnitude of density levels 5%, 0.5%, 0.037% – the last one allowing SpaRTA to have approximately the same number of trainable parameters as LoRA. LoRA has rank $r=8$ equivalent to 0.037% of trainable parameters; Head adaptation uses a very small 1.63e-4% density. Overall, the same trend as for IMDB is observed. Full FT results are very often the best, sometimes bested by SpaRTA 5% (RTE, MNLI). For a low density, where LoRA and SpaRTA use the same number of trainable parameters, SpaRTA can best LoRA (QNLI, SST2, RTE) but is overall competitive with LoRA results.
A similar set of results on the same GLUE tasks is provided for mistral-7b-it in Table 5. Again, results for the mistral-7b PT base model can be found in Table 11, see Appendix D. Full FT results can often be matched or bested by SpaRTA 5% (RTE, SST2, QQP, MNLI, MRPC, COLA). SpaRTA often gives better results than LoRA (QNLI, RTE, SST2, QQP, MNLI) for a low density (0.0048%) where both techniques have comparable numbers of parameters. Once again, SpaRTA is a competitive adaptation technique, often providing similar results to LoRA with low density.
7.4 Remarks
Results on GLUE and IMDB establish that SpaRTA is a viable adaptation technique that can be competitive with LoRA, especially for larger LMs. These results demonstrate that a simple sparsifying scheme can offer a valid adaptation technique. This opens up the possibility for more investigation in sparse adaptation as the low rank approximation of a model’s parameter changes is not the only mechanism to provide a performant adaptation method. This indicates that what matters is not necessarily the adapter structure used in PEFT techniques, but rather the number of independent parameters, that is relevant to the adaptation task at hand.
8 Exploring Model Merging
Yu et al. (2024) demonstrate that one can perform Full FT, compute $\Delta=\theta_{\text{FT}}-\theta_{\text{PT}}$, sparsify $\Delta$ to get $\tilde{\Delta}$, and obtain good performance using $\tilde{\theta}_{\text{FT}}=\theta_{\text{PT}}+\tilde{\Delta}$ as model parameters. Furthermore, they show that merging models via their sparsified $\tilde{\Delta}$-s from different tasks can also maintain the performance of the individual models. This motivates exploring model merging via SpaRTA, where no Full FT is required and sparse adapters are instead fine-tuned.
In our setup, because the heads of our SpaRTA-tuned models are fully trained, merging them into a single head would create parameter interference. We thus merge everything except the heads and add each unmerged head on top of the merged model creating a multi-head model. We need the merged model to produce multiple outputs (i.e., one for each task) per input and this is exactly what a multi-head merged model does. In the forward pass, a sequence of input tokens goes through the merged model, producing a final state representation of the tokens. This representation is shared by all tasks and is passed to each head to produce one prediction per task as output. This way all tasks are processed concurrently on the same input.
We experiment merging two models: one individually fine-tuned with SpaRTA for SST2 (classifying text sentiment) and another for COLA (classifying grammatical correctness), using gemma-2b as PT model and 99% sparsity in both cases. We merge these two models by adding the respective $\Delta_{\phi}$ to their PT model’s original parameters $\theta_{\text{PT}}$, and obtaining a two-headed model that can simultaneously classify any input sentence on both criteria in a single forward pass. Table 6 compares performance of the merged model with that of the unmerged models. This result encourages future exploration of this practical usecase for SpaRTA.
SST2 COLA
Model loss acc loss acc
SST2 0.113 96.3% - -
COLA - - 0.487 82.7%
Merged 0.118 96.7% 0.428 84.4%
Table 6: Test loss an accuracy of SST2 and COLA models fine-tuned on gemma-2b using SpaRTA with 99% sparsity. The merged model outperforms both individual models in this scenario.
9 Conclusion
As PT language models have grown in size, PEFT has become crucial for enabling fine-tuning large PT language models on limited hardware and financial resources. We have introduced SpaRTA, an approach that sharply decreases the set of trainable parameters, reducing the GPU memory used by the optimizer and speeding up training. We have demonstrated on a variety of task adaptation scenarios that our fine-tuning approach is parameter-efficient and competitive with LoRA, the current PEFT standard. Experiments with 2B and 7B parameter pre-trained models demonstrate good performance, and as per Yu et al. (2024), we expect larger models to allow for higher levels of sparsity in training; meaning that efficiency of SpaRTA should get better with larger model sizes as also suggested in Table 1.
Regarding future directions, while SpaRTA has been applied to supervised learning, it is also amenable to reinforcement learning often used for model alignment (Ouyang et al., 2022a). Furthermore, SpaRTA allows one to select which modules to sparsify and at what layers. Like LoRA, this decision affects both the number of trainable parameters and the performance of the fine-tuned model, and we plan to investigate its impact in SpaRTA. Lastly, we plan to explore model merging as discussed in Section 8; SpaRTA opens the potential for model merging without Full FT but with possibly little interference.
10 Limitations
We have demonstrated various benefits of SpaRTA, including low memory and high performance. Regarding limitations, questions remain about how to best deal with overfitting, though we have some insights. We have observed in our experiments that as we increase the sparsity level, and reduce, in turn, the number of trainable parameters, there is less overfitting. Indeed, there is a point in which both the training and validation losses converge together without significantly diverging from each other, eliminating the need for explicit overfitting mitigation techniques. Moreover, further increasing of the sparsity level beyond this point results in underfitting. Thus, we can think of our approach as a technique to improve generalization by limiting a model’s capacity to overfit to the training data. However, finding the breaking point at which this happens requires expensive experimentation. We leave as future work the investigation of how such an optimal sparsity level depends on the model, dataset sizes, and task complexity. Knowing this relation will allow us to determine in advance how much sparsity in the trainable parameters is needed for reducing the capacity of a large model to a point where it learns a new task on a relatively small amount of examples without overfitting.
Appendix A Ranks of differences in fine-tuned weight matrices
LoRA is based on the idea that the changes in the model parameters after adapting it to a new task can have a low-rank approximation. To see if this is the case, we can easily check this presumption in a few well-known fine-tuned models by comparing their weight matrices before and after fine-tuning. Thus, for every weight matrix $W_{\text{PT}}$ (i.e. a $2$-dimensional trainable tensor) in the pre-trained model, we compute the delta matrix $\Delta=W_{\text{FT}}-W_{\text{PT}}$, where $W_{\text{FT}}$ is the corresponding weight matrix after fine-tuning the model over all the original model parameters using standard supervised (or reinforcement) learning. LoRA assumes that these matrices don’t change or their differences given by $\Delta$ are low-rank.
We compute all delta weight matrices for two well-known instruction-following fine-tuned models: gemma-2b-it and mistral-7b-it, see Section 6.2. As we can see in Tables 7 and 8, the feedforward (MLP) delta $\Delta$-matrices associated with each layer of the transformer network are all full rank in both models. This is also the case for all self attention key (k) and value (v) projection matrices. The self attention query (q) and output (o) $\Delta$-matrices all show relatively small rank deficiencies: between $9$ and $2$ for the query projection matrices and between $46$ and $3$ for the output projection matrices, out of a potential maximum rank of $2,048$ for the gemma-2b-it model; and between $1,788$ and $28$ for the query projection matrices and between $246$ and $10$ for the output projection matrices, out of a potential maximum rank of $4,096$ for mistral-7b-it model. The token embedding $\Delta$-matrix for mistral-7b-it is full rank. However, it has a rank of $97$ for gemma-2b-it, being the only matrix for that model whose delta change shows a significant rank deficiency (of $1,951$). Of the two models, mistral-7b-it is the only one that does not tie its tokens embedding weights to its head; so that mistral-7b-it model’s head (untied LM head), which is trained independently from the tokens embedding, has a rank deficiency of $1$ for its delta change. Basically, we observe that the fine-tuning changes in all weight matrices of these models are full rank or very close to it, with the only exception being the changes in the tokens embedding for gemma-2b-it which are significantly low rank.
We analyzed here the changes in fine-tuned weight matrices for two popular open-weights language models. As discussed in Section 2.1, we compared instruction-following models with their original, pre-trained versions, by computing the differences (delta) in their weight matrices. Tables 7 and 8 provide a detailed breakdown of the ranks for all these delta matrices. The observed ranks suggest that constraining the (delta) changes in weight matrices to be low-rank is not essential for fine-tuning models efficiently.
Weight matrix dims Layer(s) Rank Rank deficiency
tokens embed [ 256000, 2048 ] - 97 1951
self attn q proj [ 2048, 2048 ] 0 2039 9
15 2041 7
14 2042 6
6 2044 4
1,3-5,7,8,10-13,16,17 2045 3
2,9 2046 2
self attn k proj [ 256, 2048 ] 0-17 256 0 (full rank)
self attn v proj [ 256, 2048 ] 0-17 256 0 (full rank)
self attn o proj [ 2048, 2048 ] 0 2002 46
16 2040 8
11,12,14 2042 6
6,8,10,13,15 2043 5
3-5,7,9,17 2044 4
1,2 2045 3
mlp gate proj [ 16384, 2048 ] 0-17 2048 0 (full rank)
mlp up proj [ 16384, 2048 ] 0-17 2048 0 (full rank)
mlp down proj [ 2048, 16384 ] 0-17 2048 0 (full rank)
Table 7: Rank of the change differences in gemma-2b-it model weight matrices after full fine-tuning.
Weight matrix dims Layer(s) Rank Rank deficiency
tokens embed [ 32768, 4096 ] - 4096 0 (full rank)
self attn q proj [ 4096, 4096 ] 0 2308 1788
1 3103 993
3 4023 73
31 4046 50
30 4050 46
2,4 4051 45
6 4052 44
29 4056 40
28 4057 39
11,27 4058 38
8,9 4059 37
10,19,25,26 4061 35
12,18,23 4062 34
5,7,14,17 4063 33
13,20,22 4064 32
15 4065 31
21,24 4066 30
16 4068 28
self attn k proj [ 1024, 4096 ] 0-31 1024 0 (full rank)
self attn v proj [ 1024, 4096 ] 0-31 1024 0 (full rank)
self attn o proj [ 4096, 4096 ] 0 3850 246
31 4059 37
1,30 4060 36
14 4068 28
18 4069 27
23 4070 26
2,10,26 4071 25
8,27 4072 24
3,9 4074 22
4,11,29 4075 21
6,13,22,28 4076 20
7,12 4077 19
15,16,19 4079 17
17 4080 16
5 4081 15
25 4083 13
21,24 4085 11
20 4086 10
mlp gate proj [ 14336, 4096 ] 0-31 4096 0 (full rank)
mlp up proj [ 14336, 4096 ] 0-31 4096 0 (full rank)
mlp down proj [ 4096, 14336 ] 0-31 4096 0 (full rank)
untied lm head [ 4096, 32768 ] - 4095 1
Table 8: Rank of the change differences in mistral-7b-it model weight matrices after full fine-tuning.
Appendix B Additional Dataset details
Dataset Classes Train Dev Test
COLA 2 7,551 1,000 1,043
MNLI 3 100,000 10,000 19,647
MRPC 2 3,668 408 1,725
QNLI 2 99,701 5,000 5,463
QQP 2 100,000 5,000 40,430
RTE 2 2,165 300 277
SST-2 2 66,349 1,000 872
IMDB 2 25,000 5,000 20,000
Table 9: Sequence classification datasets. Training sets limited to 100K samples. Training samples with more than 256 tokens are removed (here using mistral-7b tokenizer).
IMDB: contains a sample of “highly polar” movie reviews obtained from the online Internet Movie Database (IMDb) website. IMDb registered users provide a rating (from $1$ to $10$) with each review. The reviews are binary (positive/negative) labeled with their sentiment, defined from user ratings (Maas et al., 2011). Reviews with a rating higher or equal than $7$ are given a positive label; and a negative label if the rating is lower or equal than $4$. No reviews with ratings beyond these ranges are present in the dataset, which was constructed to have an equal number of positive and negative reviews, so guessing randomly yields 50% accuracy.
GLUE: Refer to Wang et al. (2019) which contains descriptions and details on all GLUE datasets.
Table 9 shows the actual splits used in our experiments for the IMDB and GLUE datasets, where the training data was filtered using the mistral-7b tokenizer by a threshold in the maximum token length of an network input. The only differences are for the training splits of the MRPC, QNLI, and RTE datasets.
Appendix C Training details
We adapted each (generative) pre-trained model in Section 6.2 to do sequence classification as per the examples in any of the datasets from Table 2, using different supervised fine-tuning methods, including SpaRTA. Base models were adapted by switching their vocabulary heads to randomly initialized sequence classification heads with $c$ output classification tokens. For instruction models, we re-used the vocabulary heads as described in Section 6.3.
The examples (e.g., text extracts, sentences) to be classified were converted into sequences of tokens before passing them as inputs to a model. We wrapped each example into an instruction to take advantage of instruction-following models, adding to the token length of the model inputs. We tried to keep such instructions as short as possible while still achieving a good initial performance before fine-tuning. For instance, after tokenization, the maximum token length of a training input from the SST-2 dataset was
  • 67 (without) and 87 (with instruction) for the gemma-2b-it;
  • 72 (without) and 95 (with instruction) for mistral-7b-it.
We observe here how the Gemma’s tokenizer compresses more the input than Mistral’s because of its larger vocabulary size; 256,000 (Gemma) vs. 32,767 (Mistral).
We tokenized all training examples before starting the fine-tuning and looked at the histogram of their token lengths. To avoid batches with too much padding and improve training efficiency, we dropped those training examples with a disproportionate large token length, i.e., corresponding to a tail of extreme values in the histogram. We only do this for the training data since evaluation (on the development or test data) requires much less compute and memory (no gradients need to be calculated and stored) and is performed less frequently. We will indicate in Tables 2 and 9 the final splits after filtering the training data this way, with the final number of training examples used for the fine-tuning.
With SpaRTA, we froze the token embeddings layer, made fully trainable the classification head, and randomly chose a sparse proportion of (scalar) parameters to be trained in all the other layers of a model. We demonstrated our method with varying density levels. The total average number of trainable parameters in each case was approximately: $99$M (5% density), $10$M (0.5%) and $800$k (0.037%) for the (base and instruct) Gemma 2B models; and $349$M (5% density), $35$M (0.5%) and $3.5$M (0.048%) for the (base and instruct) Mistral 7B models.
For LoRA, we factorized the changes in the query and value self-attention projection weight matrices with rank $r=8$ decomposition matrices, which were optimized while keeping all other model parameters frozen. The number of new trainable parameters introduced by the LoRA approach in the Gemma 2B and Mistral 7B models were approximately $925$k and $3.5$M, respectively. Also, we set $\alpha=16$ to scale the LoRA adapters.
We observed overfitting when training with Full parameter FT: the development loss started deteriorating after a few epochs (e.g., approximately $2$ for SST-2) while the training loss went quickly to zero. We used a combination of early stopping, dropout and weight decay to deal with overfitting. We noticed that SpaRTA is a natural regularizer: increasing sparsity resulted in less overfitting, to a point in which there was no more overfitting (e.g., this was achieved at a sparsity $s>99\%$ for all models under consideration). In general, overfitting was more noticeable with the Mistral 7B models; as expected since they are larger than the Gemma 2B models. We also noticed that larger models require higher sparsity levels to eliminate overfitting given the same training data. On the other hand, Head adapation always resulted in significant underfitting, making this fine-tuning method always underperform.
We also observed that SpaRTA achieves a speedup over Full FT for sparsity levels greater than 90%. Thus, at some sparsity level between 80% and 90%, computing gradients on a smaller amount of trainable parameters starts to offset the extra compute of adding and subtracting the deltas before and after the forward pass respectively. These extra operations require slicing the original parameter tensors over the randomly selected indices.
Appendix D Additional Experiment Results
Tables 10 and Table 11 present the complete set of results on GLUE for the PT base models gemma-2b and mistral-7b, respectively.
Model: google/gemma-2-2b QNLI RTE SST2 QQP MNLI MRPC COLA loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc Full FT 0.20 92.6 0.43 80.1 0.11 96.7 0.28 88.3 0.36 86.4 0.33 86.1 0.41 60.3 SpaRTA (5%) 0.15 94.4 0.46 80.5 0.11 96.8 0.23 90.1 0.28 89.4 0.38 83.4 0.34 64.6 SpaRTA (0.5%) 0.15 94.1 0.61 69.0 0.12 96.3 0.24 89.8 0.30 89.0 0.38 83.4 0.38 62.2 SpaRTA (0.037%) 0.29 88.3 0.69 57.8 0.11 96.2 0.30 86.9 0.47 81.1 0.58 70.3 0.52 42.1 LoRA (0.037%) 0.17 93.6 0.59 69.0 0.11 96.1 0.26 88.6 0.30 88.6 0.36 84.5 0.36 61.7 Head (1.63e-4%) 0.56 71.8 0.68 60.3 0.32 89.1 0.43 79.9 0.89 58.7 0.59 70.4 0.53 33.2
Table 10: Test loss and accuracy of model google/gemma-2-2b adapted to glue datasets with different fine-tuning methods. For training details see Table 13.
Model: mistralai/Mistral-7B-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc Full FT 0.12 95.4 0.28 89.9 0.11 96.9 0.23 90.1 0.25 91.1 0.34 85.9 0.32 66.8 SpaRTA (5%) 0.12 95.5 0.34 86.3 0.11 96.9 0.23 90.3 0.25 90.9 0.34 87.6 0.37 64.3 SpaRTA (0.5%) 0.12 95.9 0.35 86.3 0.11 96.7 0.22 90.5 0.24 91.2 0.32 88.8 0.34 68.2 SpaRTA (0.048%) 0.13 95.3 0.36 86.3 0.11 96.8 0.24 89.7 0.26 90.3 0.38 86.3 0.34 66.9 LoRA (0.048%) 0.12 95.6 0.30 88.1 0.11 96.7 0.23 90.0 0.25 91.0 0.31 87.3 0.32 67.8 Head (1.15e-4%) 0.39 82.8 0.63 70.4 0.16 94.8 0.33 85.8 0.63 74.0 0.53 73.8 0.48 45.5
Table 11: Test loss and accuracy of model mistralai/Mistral-7B-v0.3 adapted to glue datasets with different fine-tuning methods. For training details see Table 13.
Appendix E Training Hyper-Parameters
The hyper-parameters used for investigating SpaRTA and other adaptation methods on IMDB and GLUE are summarized next.
E.1 IMDB
The sets of best parameters used in training the various adaptation methods for each model are given in Table 12. To improve training efficiency, we excluded training examples exceeding 384 tokens in length, resulting in the use of only 19,306 examples for gemma-2b, 18,744 examples for gemma-2b-it, 18,373 examples for mistral-7b, and 17,672 examples for mistral-7b-it for training. That is 77%, 75%, 73%, and 71% of the original training data of 25,000 examples, respectively.
Parameter gemma-2b gemma-2b-it mistral-7b mistral-7b-it Full FT batch size 32 32 36 36 num epochs 2 2 $\ast$ $\ast$ learning rate 1e-5 1e-5 3e-6 3e-6 max grad norm 10 50 120 120 dropout 0.1 0.1 0.15 0.15 weight decay 0.1 0.1 0.01 0.01 SpaRTA batch size 40 40 16 16 num epochs. 2 2 $\ast$ $\ast$ $d=$ 5% learning rate 1.5e-5 8e-6 2e-6 2e-6 $d=$ 0.5% learning rate 1e-4 5e-5 3e-6 3e-6 $d=$ 0.05% learning rate 6e-5 6e-5 6e-5 6e-5 LoRA batch size 40 40 20 20 num epochs 3 3 3 3 learning rate 2e-4 2e-4 5e-6 5e-6 max grad norm 15 15 - - dropout 0.1 0.1 0.1 0.1 $r$ 8 8 8 8 $\alpha$ 16 16 16 16 Head batch size 40 40 16 16 num epochs 4 4 3 3 learning rate 2e-4 2e-4 1e-4 1e-4
Table 12: (IMDB) Training parameters used with each fine-tuning method and model in Table 3. $\ast$ in number of epochs indicates early stopping was used. For SpaRTA, parameters for density 5%, 0.5% 0,05% are reported.
E.2 GLUE
For GLUE, given the number of tasks (7) and the number of methods tried (6), reporting each best set of training parameters would be just prohibitive. Instead, the trainings of these methods was done using a simple grid search over some of the hyper-parameters. The range on these hyper-parameters are reported in Table 13. Similarly to IMDB, any sample longer than 256 tokens (which is dependnent on the tokenizer used) is discarded from the dataset to avoid having few minibatch with one very long sample compared to the rest, as seen in Table 2.
Parameter Grid Search Value Range
Full FT batch size [32, 24, 8] depending on task and model to fit in GPU memory
learning rate [1e-3, 5e-4, 2e-4, 1e-4, 5e-5, 1e-5, 5e-6]
dropout [0.1]
weight decay [0.1]
SpaRTA batch size [same as for Full FT]
learning rate [same as for Full FT]
weight decay [0.1]
LoRA batch size [same as for Full FT]
learning rate [same as for Full FT]
$r$ [8]
$\alpha$ [16]
dropout [0.1]
weight decay [0.1]
Head batch size [same as for Full FT]
learning rate [same as for Full FT]
dropout [0.1]
weight decay [0.0]
Table 13: (GLUE) Training parameters and their grid search value range to get results for each adaptation method for each model in Table 4, Table 5, Table 10, and Table 11. Best set of parameters were chosen from best validation loss results over 2 epochs of training.

### v2

@@
 1 Introduction
-It has become common practice to train application-ready language models in two phases (Radford et al., 2018; Kenton and Toutanova, 2019): first, the model is pre-trained on a very large and general corpus of (unlabeled) text; then further trained (or fine-tuned) on a smaller specific set of examples demonstrating the intended behavior for a particular application, such as instruction following (Ouyang et al., 2022b).
+It has become common practice to train application-ready language models in two phases (Radford et al., 2018; Kenton and Toutanova, 2019): first, the model is pre-trained on a very large and general corpus of (unlabeled) text; then further trained (or fine-tuned) on a smaller specific set of examples demonstrating the intended behavior for a particular application, such as instruction following (Ouyang et al., 2022).
 Overall, supervised fine-tuning (SFT) requires less computational resources than pre-training (PT) due to the significantly smaller size of the training set as well as the typical use of early stopping to deal with overfitting.
@@
 Generally, one samples $m$ parameters from the pre-trained model $\theta_{\text{PT}}$, stores the indices of these parameters in $\phi$, and uses adapter $\Delta_{\phi}$ to fine-tune $\theta_{\text{PT}}$.
-To investigate the performance of SpaRTA, we build adapters with different sparsity levels and test on the GLUE classification tasks (Wang et al., 2019) as well as the IMDB dataset (Maas et al., 2011).
-SpaRTA is compared to a fully fine-tuned model (Full FT), and PEFT approaches including LoRA.
-SpaRTA is quite competitive compared to LoRA given that it only modifies a small sparse number of model parameters.
+To investigate the performance of SpaRTA, we build adapters with different sparsity levels and evaluate them on a wide range of natural language understanding benchmarks.
+SpaRTA is compared to other PEFT approaches including LoRA and DoRA Liu et al.
+(2024), and found to be quite competitive compared to these methods given that it only modifies a small sparse number of model parameters.
 2 Motivation
 Yu et al.
-(2024) look at the differences between a language model’s parameters before and after fine-tuning, and demonstrate empirically that it is possible to randomly drop (i.e., set to zero) up to 99% of these parameter changes without significantly affecting the model performance.
-The sparsity in the parameter changes represented by the tensor $\Delta$ is known as $\Delta$-sparsity.
+(2024) look at the differences between a language model’s parameters before and after fine-tuning, and demonstrate empirically that it is possible to randomly drop (i.e., set to zero) up to 99% of these parameter changes, represented by $\Delta$, without significantly affecting model performance.
 This motivates our approach, SpaRTA, which produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters: SpaRTA randomly selects the (scalar) parameters to train and freezes the remaining parameters (i.e., setting the corresponding $\Delta$ values to zero).
-This parameter sparsity also helps in reducing overfitting, as pre-trained models typically have enough capacity to learn the often limited amount of (labeled) data used for fine-tuning.
-There is no guarantee of $\Delta$-sparsity in LoRA, but this is a desired property since it reduces parameter interference (Yadav et al., 2023) when merging fine-tuned, task-specific models into a single one that can perform all tasks simultaneously.
-One byproduct of $\Delta$-sparsity is sparse $\Delta$ updates which reduce gradient computations during training, ensure fast merge back of the sparse $\Delta$ into the model for inference, and ultimately same inference cost as the original model (like LoRA).
+This $\Delta$-sparsity also helps in reducing overfitting, as pre-trained models typically have more than enough capacity to learn the often limited amount of (labeled) data used for fine-tuning.
+There is no guarantee of $\Delta$-sparsity in LoRA, but this is a desired property since it reduces parameter interference (Yadav et al., 2023) when merging fine-tuned models.
+SpaRTA, like LoRA, reduces the number of gradient computations during training (when compared with full Fine-Tuning), and ultimately has the same inference cost as the original model (after merging back the sparse $\Delta$ into the model).
 The properties of low memory, low computation needs in training, identical inference time as the original model are all quite desirable for an adaptation technique.
 SpaRTA has all these properties plus the unique added benefit of producing only sparse changes in a small number of the parameters of the original model that can be beneficial for merging multiple SpaRTA adapters.
-2.1 Is low-rank necessary for $\Delta$?
-In Appendix A, we show that the weight matrix changes (i.e., $\Delta$) during full parameter fine-tuning are, in fact, not generally low-rank for capable models such as gemma-2b and mistral-7b.
-This indicates that LoRA works, not because of the low-rank constraint particularly, but rather due to the capacity reduction achieved since fine-tuning with LoRA is done with limited training data.
+2.1 Is low-rank adaption necessary?
+In Appendix A, we show that the changes (i.e., $\Delta$) in weight matrix during full parameter fine-tuning are, in fact, not generally low-rank for capable models such as gemma-2b and mistral-7b.
+This indicates that LoRA works, not particularly because of its low-rank constraint, but rather due to the capacity reduction achieved since LoRA fine-tuning with is typically done on limited training data.
 Such insight also hints that any constraint which reduces the capacity of the original model could perform competitively, motivating our approach that selects a small number of parameters of the original model to be updated during training.
@@
 The last few years has seen many advances in fine-tuning Large Language Models (LLMs).
-Perhaps the most well-known method (also most used in practice) is LoRA Hu et al.
-(2022), which has spurred many variants including DoRA – which adapts only the directions of pre-trained weights Liu et al.
-(2024), VeRA – which shares low-rank matrices across layers Kopiczko et al.
-(2024), AdaLoRA – which adaptively allocates parameter budgets among weight matrices according to their importance Zhang et al.
-(2023), and SoRA – which dynamically adjusts the intrinsic rank during the adaptation Ding et al.
+Perhaps the most well-known and used method in practice is LoRA Hu et al.
+(2022), which has spurred many variants including: DoRA, which adapts only the directions of pre-trained weights Liu et al.
+(2024); VeRA, which shares low-rank matrices across layers Kopiczko et al.
+(2024); AdaLoRA, which adaptively allocates parameter budgets among weight matrices according to their importance Zhang et al.
+(2023); and SoRA, which dynamically adjusts the intrinsic rank during adaptation Ding et al.
 (2023).
-This set of methods each have a different structure to the $\Delta$ being optimized with the commonality being the training of some function of low-rank matrices.
-For example, further expanding on VeRA Kopiczko et al.
-(2024), the authors propose to learn $\Delta=DBEA$ where $A$ and $B$ are random low-rank matrices and $D$ and $E$ are diagonal matrices to be trained.
+Each of these methods has a different structure to the $\Delta$ being optimized, with the commonality being the training of some function of low-rank matrices.
+AdaLoRa and SoRA are also dynamic methods, which adjust the function during adaptation.
 Beyond adding structured parameters is the concept of fine-tuning a small subset of the total parameters of the model, i.e., sparse fine-tuning, where one must first decide which subset of parameters to fine-tune (similar to deciding which parameters in each layer to add adapters to in LoRA).
@@
 (2022) first fine-tune on all parameters, select the parameters that change the most, and then fine-tune again from scratch on the selected parameters.
-In contrast to these, our proposed SpaRTA method produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters chosen completely at random.
+Deng et al.
+(2025) proposes improvements via a modified pruning strategy to the random dropping and rescaling method of Yu et al.
+(2024) which motivates this work.
+In contrast to these more complex, sometimes dynamic approaches, our proposed SpaRTA method produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters chosen completely at random.
 Yet another related direction is that of compression which results in sparse models; algorithms in this genre take a dense fine-tuned model with the goal of compressing it while maintaining performance.
@@
 Suppose the vectorized parameters of the pre-trained language model are $\theta_{\text{PT}}\in\mathbb{R}^{n}$ where $n$ is the number of parameters, and full parameter fine-tuning (FT) of the model is performed with a labeled dataset characterizing a task.
-FT typically updates $\theta_{\text{PT}}$ using a stochastic first-order gradient-based optimization algorithm (e.g., Adam from Kingma and Ba (2015)) to maximize the conditional probability of the labels in the training dataset under the model.
+FT typically updates $\theta_{\text{PT}}$ using a stochastic first-order gradient-based optimization algorithm, e.g., Adam from Kingma and Ba (2015), to maximize the conditional probability of the labels in the training dataset under the model.
 The FT model is then given by $\theta_{\text{FT}}=\theta_{\text{PT}}+\Delta_{\text{FT}}$ where $\Delta_{\text{FT}}\in\mathbb{R}^{n}$.
 This has two drawbacks in terms of memory efficiency.
-First, the optimizer state can be large, e.g., the state of the Adam optimizer is $4$ times as large as the parameter space as it includes current parameter values, a moving average of gradients (per parameter), etc.
+First, the optimizer state can be large, e.g., the state of the Adam optimizer is $4$ times as large as the parameter space as it includes current parameter values, their gradients as well as per parameter statistics of those gradients.
 Second, storing a new FT model takes as much memory as the PT model which could be an issue when many task-specific models are requested.
@@
 The model parameters $\theta\in\mathbb{R}^{n}$ are partitioned into trainable $\theta_{\text{T}}\in\mathbb{R}^{m}$ and frozen $\theta_{\text{F}}\in\mathbb{R}^{n-m}$ ones, with $m\ll n$ being the number of parameters selected.
-This allows our approach to have a drastically lower memory footprint than FT (similar to LoRA) by reducing the size of the optimizer state as well as faster training by reducing the number of gradients to compute.
-In our SpaRTA’s implementation, we introduce: (i) non-trainable indices $\phi\in\mathbb{R}^{m}$ containing the indices of the randomly selected elements in $\theta_{\text{PT}}$, and (ii) trainable parameters $\Delta_{\phi}\in\mathbb{R}^{m}$ representing the subset of $\Delta$ that SpaRTA learns at indices $\phi$.
+This allows our approach to have a drastically lower memory footprint than FT by reducing the size of the optimizer state as well as faster training by reducing the number of gradients to compute.
+In our SpaRTA’s implementation [^1] https://github.com/IBM/SpaRTA, we introduce: (i) non-trainable indices $\phi\in\mathbb{R}^{m}$ containing the indices of the randomly selected elements in $\theta_{\text{PT}}$, and (ii) trainable parameters $\Delta_{\phi}\in\mathbb{R}^{m}$ representing the subset of $\Delta$ that SpaRTA learns at indices $\phi$.
 The pseudocode for SpaRTA is given in Algorithm 1.
@@
 Input: Task labeled dataset $\mathcal{D}$
-Sample $\phi$: Indices of parameters to be optimized
+Sample $\phi$: Indices of (scalar) parameters to be optimized
 Initialize: $\Delta_{\phi}=0$
@@
 Whereas optimizing LoRA requires computing gradients with respect to the $A$ and $B$’s used to compute $\Delta$ (which constitute additional parameters, independent from $\theta_{\text{PT}}$), SpaRTA requires computing gradients with respect to $\Delta_{\phi}$, the changes over a subset of parameters chosen from $\theta_{\text{PT}}$ and indexed by $\phi$.
-This nuance is why LoRA can trained with out-of-the-box optimizers, but SpaRTA requires a different implementation.
-For example, Step 4 in Algorithm 1 must take into account the computational graph so as to not interfere with the forward pass to be computed at the following iteration.
-Specifically, this unmerge operation must not be recorded in the computational graph.
-Generating the index set $\phi$ is done as follows.
-For each $i\in\{1,\ldots,n\}$, SpaRTA samples $x\sim\mbox{Bernoulli}(m/n)$ and includes $i\in\phi$ if $x=1$.
-Hence, SpaRTA uses $m$ trainable parameters in expectation.
-Step 5 also depends on the specific optimizer and settings being used; SpaRTA uses Adam in our implementation.
+Generating the index set $\phi$ in SpaRTA is done by sampling from a Bernoulli independently and including each scalar parameter in $\phi$ with probability $m/n$.
+Hence, SpaRTA uses $m$ trainable parameters out of $n$ in expectation.
 Finally, inference is performed, similarly to LoRA, by merging the fine-tuned $\Delta$ into $\theta_{\text{PT}}$ and then making the necessary forward passes.
@@
 In contrast, SpaRTA only optimizes $\Delta_{\phi}$, requiring a total of $m\,(4\times 16+2\times 16)+n\times 16$ bits of memory: (i) $m\,(4\times 16)$ bits for the statistics needed by Adam to optimize $\Delta_{\phi}$ of size $m$, (ii) $m\,(2\times 16)$ bits for $\phi$ with the indices identifying the model parameters associated with $\Delta_{\phi}$, and (iii) the PT model parameters ($n\times 16$ bits).
-Memory savings appear if and only if
+Memory savings appear then if and only if
 $m\,(4\!\times\!16+2\!\times\!16)+n\times 16$ $<4n\!\times\!16,$ (1)
@@
 For instance, using SpaRTA on a model with sparsity $s=$ 80%, 90%, 95%, and 99% requires 45%, 60%, 67.5% and 73.5% less memory than full parameter FT, respectively, see Table 1.
+Additional savings are possible regarding storing indices given a fixed random number generator; in this case, the random path of selected parameters to train can be derived given a fixed seed.
 Model storage Full FT SpaRTA $n$ (on disk) 0% 50% 80% 90% 95% 99% 2B 4 16 16 8.8 6.4 5.2 4.2 7B 14 56 56 30.8 22.4 18.2 14.8
@@
 6 Experimental Setup
-We next detail the framework for conducting the experiments in Section 7.
-Motivation is first given for the tasks followed by a description of models to be used and how these models are adapted (with manipulations) for the desired tasks.
+We next detail our experimental framework. Motivation is first given for the tasks followed by a description of models to be used and how these models are adapted (with manipulations) to the tasks.
 6.1 Tasks
 The main experimental focus is on Natural Language Understanding (NLU) tasks, specifically sequence classification, which involves classifying natural language sequences into a given number of classes.
-NLU tasks are easier to evaluate than Natural Language Generation (NLG) tasks, as one can use simple metrics (Accuracy, F1, Recall, Precision, etc.) that can avoid the inherent ambiguity of evaluating more general generative tasks.
+NLU tasks are easier to evaluate than Natural Language Generation (NLG) tasks, as one can use simple metrics (Accuracy, F1-scores, etc.) that can avoid the inherent ambiguity of evaluating more general generative tasks.
 While SpaRTA is also applicable to NLG tasks, they are not used in the following demonstrations due to the challenges associated with their evaluation, typically requiring human judgments as the gold standard measure of performance.
@@
 6.2 Language Models
-Starting with available open-weight language models, the goal is to adapt them to perform sequence classification.
+Starting with available open-weight language models, the goal is to adapt them to perform a selected NLU task.
 Two types of trained models are used: base and instruction-tuned models, where the latter have additionally been trained to follow users’ intent when prompted, such as answering questions or responding to instructions.
-Specifically, we consider the following language models: gemma-2b and gemma-2b-it from the Gemma (Team et al., 2024) family, and mistral-7b and mistral-7b-it from the Mistral [^2] We use v0.3 for both models.
-(Jiang et al., 2023) family, developed by Google and Mistral AI respectively.
+Specifically, we consider the following language models: gemma-2b and gemma-2b-it from the Gemma family (Team et al., 2024), and mistral-7b and mistral-7b-it from the Mistral [^3] We use v0.3 for both models.
+family (Jiang et al., 2023).
 The "it" suffix refers to instruction-tuned models.
 They are of particular interest for our experiments as they will show results for models with two different numbers of parameters.
-All models are readily available text-to-text, decoder-only transformer models with open weights, downloadable from Hugging Face [^3] https://huggingface.co/(HF).
+All models are readily available text-to-text, decoder-only transformer models with open weights, downloadable from Hugging Face [^4] https://huggingface.co/.
 Note that when using an instruction-following (it) model, inputs are formatted according to the conventions established when training the model on instructions.
 6.3 Task Adaptation
-When using a base model for sequence classification, the raw sequences to be classified are input directly into the model.
+When using a base model for a sequence classification task, the raw text to be classified is input directly into the model as a sequence of tokens (with some special token if necessary to deal with structured inputs).
 However, when using an instruction model, these sequences are first wrapped into a classification-specific instruction.
@@
 Efficiently adapting this model for sequence classification requires the swap of this head for a sequence classification head, which uses only the final hidden state of the last token in the input sequence $h\in\mathbb{R}^{d}$ to do the classification.
-This reduces the parameter size of the head, which is just a linear layer with no bias applied to a token’s final hidden state, from a weight matrix $W\in\mathbb{R}^{v\times d}$ to $W\in\mathbb{R}^{c\times d}$, where $d$ is the dimension of the model’s hidden states (e.g., 2,048 and 4,096 for the gemma-2b and mistral-7b models respectively), $v$ is the number of tokens in the vocabulary (e.g., 256,000 for gemma-2b or 32,768 for mistral-7b), and $c$ is the number of classes in the individual task (e.g.
-$2$ to $3$ in the experiments that follow).
-With this small change, the model outputs classification probabilities for each whole input sequence through
+This reduces the parameter size of the head, which is just a linear layer applied to $h$, from a weight matrix $W\in\mathbb{R}^{v\times d}$ to $W\in\mathbb{R}^{c\times d}$, where $d$ is the dimension of the model’s hidden states (e.g., 2,048 and 4,096 for the gemma-2b and mistral-7b models respectively), $v$ is the number of tokens in the vocabulary (e.g., 256,000 for gemma-2b or 32,768 for mistral-7b), and $c$ is the number of classes (e.g.
+$2$ to $4$ in the experiments that follow).
+With this small change, the model outputs classification probabilities for each input sequence through
 $p=\operatorname*{softmax}(h\,W^{T})\in\mathbb{R}^{c}.$ (2)
-The classification head of our base pre-trained models (i.e., gemma-2b and mistral-7b) is initialized with random weights, sampled from a normal distribution $\mathcal{N}(0,\sigma^{2})$, with $\sigma$ chosen such that the model at initialization predicts either class with equal probability.
-This is achieved with $\sigma=0.06/\sqrt{d}$.
-With this, the loss starts at approximately $-\log(1/2)\approx 0.693$ and the accuracy at $\approx 50\%$ for a balanced evaluation dataset.
-Regarding instruction-tuned models (i.e., gemma-2b-it and mistral-7b-it), the original vocabulary weights are rather reused when initializing their classification heads.
+The classification heads of our base pre-trained models (i.e., gemma-2b and mistral-7b) are initialized with random weights.
+The weights of the the original vocabulary heads of instruction-tuned models (i.e., gemma-2b-it and mistral-7b-it) are rather reused when initializing their classification heads.
 To do so requires to first identify the tokens in the vocabulary that the model is expected to use for classification following the instruction.
-For example, these could be the tokens associated with a “Yes” or “No” response by the model.
+For example, these could be the tokens associated with a “Yes” or “No” response.
 The embeddings in the original model (head) associated with those classification tokens are extracted and used to initialize the classification head.
@@
 7 Experimental Results
-Efficacy of SpaRTA is demonstrated empirically by comparing it to the following baselines:
-  • Full parameter Fine-Tuning (Full FT): all model parameters are optimized.
-  • LoRA: Only the additional matrices representing low-rank adapters are optimized.
-  • Head: Only the classification head is optimized; all other parameters remain frozen.
-This enables the exploration of the performance of SpaRTA on a range of sparsity levels, varying the number of trainable parameters.
-A complete set of results can be found in Appendix D accompanied by a detailed description of the training setup in Appendices C and E.
-7.1 Datasets
-Each sequence classification task considered in our experiments is given by a dataset of labeled examples.
-Table 2 summarizes these datasets and their splits for training, development, and testing for both the GLUE classification tasks (Wang et al., 2019) and IMDB (Maas et al., 2011).
-See Appendix B for more detailed descriptions.
+Efficacy of SpaRTA is demonstrated empirically by comparing it to two baselines, LoRA (only the additional matrices representing low-rank adapters are optimized) and DoRA (magnitude is optimized in addition to low-rank adapters).
+We also explore the performance of SpaRTA on a range of sparsity levels, varying the number of trainable parameters.
 Dataset Classes Train Dev Test
+IMDB 2 25,000 5,000 20,000
 COLA 2 7,551 1,000 1,043
 MNLI 3 100,000 10,000 19,647
-MRPC 2 3,668 408 1725
+MRPC 2 3,668 408 1,725
 QNLI 2 99,725 5,000 5,463
@@
 SST-2 2 66,349 1,000 872
-IMDB 2 25,000 5,000 20,000
+BoolQ 2 9,427 1,270 2,000
+MMLU 4 99,842 1,531 14,042
 Table 2: Sequence classification datasets.
@@
 Training samples with > 256 tokens are removed (here using gemma-2b tokenizer, with mistral-7b tokenizer in Appendix B).
+We consider several NLU benchmarks, including IMDB (Maas et al., 2011), GLUE (Wang et al., 2019), BoolQ (Clark et al., 2019) and MMLU (Hendrycks et al., 2020).
+See Appendix B for detailed descriptions.
+Table 2 summarizes these datasets and our splits for training, development, and testing.
+A detailed description of the training setup can be found in Appendices C and E.
+All results are averaged over 3 random seeds.
 7.1 IMDB
-gemma-2b gemma-2b-it mistral-7b mistral-7b-it Adaptation density (%) loss acc.
+Sparsification targets Loss Accuracy
+Wq, Wv (like LoRA) 0.117 96.3%
+Wv, Wo 0.109 96.7%
+Wq, Wk, Wv 0.123 95.9%
+Wq, Wk, Wo 0.112 96.0%
+Wq, Wq, Wv, Wo 0.110 96.5%
+MLP 0.150 94.7%
+Wq, MLP 0.150 94.5%
+Wk, MLP 0.150 95.1%
+Wv, MLP 0.145 95.2%
+Wo, MLP 0.140 95.6%
+Wq, Wk, MLP 0.150 94.6%
+Wv, Wo, MLP 0.136 95.8%
+W, MLP, norm 0.140 95.4%
+Table 3: Full table of loss and accuracy on the SST-2 task after applying our SpaRTA approach to different types of parameters in the Gemma 2B model, given that the same number of trainable parameters are selected.
+Results are averaged across 10 random seeds
+gemma-2b gemma-2b-it mistral-7b mistral-7b-it method density (%) loss acc.
 loss acc.
@@
 loss acc.
-Full FT 100% 0.092 96.9% 0.107 96.3% 0.080 97.4% 0.080 97.3% SpaRTA 5% 0.096 96.7% 0.105 96.3% 0.084 97.2% 0.081 97.5% SpaRTA 0.5% 0.096 96.7% 0.103 96.4% 0.087 96.9% 0.087 97.0% SpaRTA 0.05% 0.106 96.3% 0.114 96.2% 0.080 97.3% 0.076 97.4% LoRA $\approx$0.050.101 96.5% 0.113 96.2% 0.086 97.1% 0.081 97.1% Head $\approx$2e-40.165 93.6% 0.194 92.7% 0.139 94.9% 0.117 95.8%
-Table 3: Test loss and accuracy of models adapted to the IMDB dataset with different fine-tuning methods. For training details see Table 12.
-Table 3 shows the results for the IMDB classification task where each model is asked to rate a review as positive or negative.
-Each model is fine-tuned using adaptation techniques including: (i) Full FT where all model parameters are updated (100% density); (ii) SpaRTA for different orders of magnitude of density levels 5%, 0.5%, $\approx$0.05– the last one allowing for SpaRTA to have close to the same number of trainable parameters as for LoRA; (iii) LoRA with rank $r=8$ equivalent to $\approx 0.05\%$ of trainable parameters compared to the model full parameter size; (iv) Head adaptation where only the classification head is updated ($\approx$2e-4density).
-Model: gemma-2-2b-it QNLI RTE SST2 QQP MNLI MRPC COLA loss acc.
+PT (zero-shot) 0.425 92.0% 0.435 86.1% Full FT 100% 0.092 96.9% 0.107 96.3% 0.080 97.4% 0.080 97.3% SpaRTA 5% 0.096 96.7% 0.105 96.3% 0.084 97.2% 0.081 97.5% SpaRTA 0.5% 0.096 96.7% 0.103 96.4% 0.087 96.9% 0.087 97.0% SpaRTA 0.05% 0.106 96.3% 0.114 96.2% 0.080 97.3% 0.076 97.4% LoRA $\approx$0.050.101 96.5% 0.113 96.2% 0.086 97.1% 0.081 97.1%
+Table 4: Test loss and accuracy of models adapted to the IMDB dataset with different fine-tuning methods.
+We also report zero-shot performance of instruction following PT models.
+For training details see Table 16.
+Model: gemma-2b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc.
 loss acc.
@@
 loss acc.
-loss mcc Full FT (100%) 0.46 81.2 0.33 87.4 0.18 93.3 0.36 85.5 0.26 88.9 0.12 96.1 0.36 63.6 SpaRTA (5%) 0.45 79.8 0.33 87.6 0.18 93.0 0.40 81.3 0.24 89.9 0.21 95.6 0.41 57.2 SpaRTA (0.5%) 0.55 72.9 0.37 86.1 0.25 89.8 0.34 85.4 0.24 89.9 0.13 95.6 0.39 56.9 SpaRTA (0.037%) 0.45 79.8 0.38 85.5 0.18 93.0 0.40 82.3 0.27 88.6 0.12 95.8 0.45 56.5 LoRA (0.037%) 0.46 78.3 0.33 87.5 0.24 90.4 0.36 84.8 0.26 89.0 0.13 95.5 0.39 59.9 Head (1.63e-4%) 2.60 47.3 0.73 67.4 0.96 37.1 0.55 74.2 0.66 64.9 0.55 84.2 0.62 18.0
-Table 4: Test loss and accuracy of model google/gemma-2-2b-it adapted to glue datasets with different fine-tuning methods. For training details see Table 13.
-In Table 3, the results for adaptation techniques (rows) are sorted in order of descending density as to show the impact of decreasing the number of trainable parameters on the overall performance.
-For gemma-2b and gemma-2b-it, Full FT adaptation gives strong test loss and accuracy numbers, often the best available – this is expected since all model parameters are fine-tuned.
+loss mcc loss acc.
+avg.
+PT (zero-shot) 1.34 40.5 1.77 43.7 0.75 60.7 1.41 48.3 2.91 30.9 0.91 64.6 0.85 0.7 0.88 68.5 44.7 SpaRTA (5%) ALL 0.17 93.7 0.54 81.0 0.14 95.0 0.26 89.0 0.33 87.2 0.35 85.3 0.39 56.2 0.36 84.4 84.0 SpaRTA (0.5%) 0.16 94.0 0.43 80.7 0.15 94.7 0.24 89.7 0.32 87.8 0.34 86.5 0.45 54.9 0.37 84.7 84.1 SpaRTA (0.037%) 0.18 93.1 0.48 76.9 0.16 94.5 0.26 88.8 0.35 86.4 0.36 84.5 0.45 54.8 0.40 83.6 82.8 LoRA (0.037%) Q,V 0.18 93.4 0.44 78.7 0.14 95.1 0.26 89.1 0.33 87.8 0.33 85.4 0.41 55.0 0.36 84.0 83.6 DoRA (0.037%) 0.18 93.4 0.45 78.0 0.15 95.6 0.26 89.0 0.33 87.8 0.33 85.5 0.41 57.1 0.36 84.4 83.8 SpaRTA (0.037%) 0.19 92.8 0.46 77.6 0.16 94.8 0.26 88.6 0.36 86.3 0.36 84.2 0.45 54.5 0.39 83.4 82.8 SpaRTA (0.037%) O,V 0.17 93.5 0.44 80.7 0.15 94.9 0.26 88.9 0.34 87.2 0.35 85.7 0.42 55.9 0.37 85.0 84.0
+Table 5: Test loss and accuracy of model gemma-2b-it adapted to GLUE and BoolQ datasets with different fine-tuning methods. For training details see Appendix E.2.
+Table 4 shows the results for IMDB, where each model is asked to classify a review as positive or negative.
+Each model is fine-tuned using the following adaptation methods: (i) Full parameter fine-tuning (Full FT) where all model parameters are optimized; (ii) SpaRTA for different density levels 5%, 0.5%, 0.05%, with the last allowing SpaRTA to have approximately the same number of trainable parameters as LoRA; (iii) LoRA with rank $r=8$ equivalent to about $0.05\%$ of trainable parameters compared to the model full parameter size.
+In Table 4, results for adaptation methods (rows) are sorted in order of descending density as to show the impact of decreasing the number of trainable parameters on the overall performance.
+For gemma-2b and gemma-2b-it, Full FT adaptation gives (practically always) the best test loss and accuracy numbers.
+This is expected since all model parameters are fine-tuned.
 SpaRTA results at 5% density are close to Full FT, even matching them for gemma-2b-it.
 As the density decreases by orders of magnitude, the results slowly degrade.
-For $\approx$0.05%, gemma models’ performances are slightly behind or matching performance from LoRA.
-Head adaptation provides the worse results as one would expect since it has so few parameters to work with.
-For mistral models, the trend is similar with Full FT adaptation showing strong loss and accuracy numbers.
+For 0.05% density, gemma models’ performances match or slightly lag performance from LoRA.
+For mistral models, the trend is similar with Full FT showing best (or close to best) loss and accuracy numbers.
 SpaRTA performs well, matching and even improving upon Full FT with a density of 0.05%.
 SpaRTA even improves over LoRA for both mistral-7b and mistral-7b-it at this low density.
-Once again, Head adaptation gives the worst results.
-Overall these results are encouraging and show that SpaRTA is competitive and provides similar performance to LoRA for low densities.
-7.3 GLUE
-Model: Mistral-7B-Instruct-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA loss acc.
+Overall these results are encouraging; SpaRTA is competitive, provides similar performance to LoRA for low densities, and does not degrade the performance of full parameter fine-tuning.
+7.2 Which parameters should SpaRTA target?
+So far, SpaRTA has chosen trainable parameters from any layer of the transformer with equal probability. In contrast, LoRA concentrates adaption on the key and value self-attention weight matrices.
+Now, we investigate whether it is better to concentrate the selection of sparse, trainable parameters on a few parameters or equally across all.
+Specifically, given a budget for the number of parameters that can be trained, which parameters should we target for sparsification (while freezing the remaining) to achieve best task performance when adapting the sparsified parameters within the targeted types?
+We conduct an ablation study to answer this question.
+The Gemma 2B model is adapted to the SST-2 task using our SpaRTA approach targeting different combinations of parameters.
+We set a budget of 1.25M trainable parameters within the Gemma 2B model so that sparsity $s=99.95\%$.
+Results of these experiments are shown in Table 3.
+Concentrating the selection of trainable parameters through sparsification in the self-attention value (Wv) or/and output (Wo) weight matrices yields the best performance under the given budget.
+In contrast, distributing our sparsification across different combinations of weight types may lead to significantly lower performance.
+Note that concentrating our sparsification over targeted parameters, although proven beneficial in terms of single-task adaption performance, can decrease performance when merging since it increases the chance of parameter interference.
+7.3 GLUE and BoolQ
+Model: Mistral-7B-Instruct-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc.
 loss acc.
@@
 loss acc.
-loss mcc Full FT (100%) 0.32 91.7 0.26 91.3 0.14 95.7 0.33 88.6 0.23 90.3 0.11 96.7 0.32 70.5 SpaRTA (5%) 0.30 89.9 0.23 91.4 0.12 95.8 0.28 88.9 0.23 90.5 0.10 96.7 0.30 72.1 SpaRTA (0.5%) 0.29 89.5 0.23 91.6 0.13 95.2 0.29 88.6 0.25 89.2 0.10 96.6 0.32 70.3 SpaRTA (0.048%) 0.27 89.9 0.24 91.1 0.12 95.9 0.31 87.6 0.25 89.3 0.10 96.6 0.32 69.9 LoRA (0.048%) 0.32 88.8 0.26 90.4 0.12 95.6 0.32 87.2 0.27 88.5 0.11 96.9 0.30 71.3 Head (1.15e-4%) 0.31 86.3 0.35 87.0 0.23 90.4 0.35 84.5 0.32 85.5 0.15 93.5 0.36 62.1
-Table 5: Test loss and accuracy of the mistralai/Mistral-7B-Instruct-v0.3 model adapted to glue datasets with different fine-tuning methods. For training details see Table 13.
-Table 4 compares the results of gemma-2b-it on 7 of the GLUE tasks for the same adaptation methods as in Table 3.
+loss mcc loss acc.
+avg.
+PT (zero-shot) 2.38 23.4 2.54 33.6 0.66 66.3 0.85 57.5 2.42 35.5 0.70 70.0 0.72 25.2 0.39 85.2 49.6 SpaRTA (5%) ALL 0.11 96.0 0.26 89.8 0.12 95.9 0.22 90.7 0.23 91.6 0.32 87.3 0.36 62.6 0.24 90.8 88.1 SpaRTA (0.5%) 0.11 95.8 0.30 90.5 0.12 96.0 0.23 90.4 0.24 91.2 0.32 88.0 0.35 67.9 0.24 90.5 88.8 SpaRTA (0.048%) 0.12 95.7 0.31 89.2 0.12 96.1 0.23 90.5 0.24 91.1 0.31 88.2 0.34 67.3 0.26 89.5 88.4 LoRA (0.048%) Q,V 0.12 95.6 0.24 91.2 0.12 96.0 0.23 90.2 0.24 91.0 0.29 89.3 0.32 69.9 0.24 90.5 89.2 DoRA (0.048%) 0.13 95.0 0.27 90.9 0.11 96.4 0.23 90.2 0.25 90.9 0.28 88.7 0.32 68.2 0.24 90.5 88.8 SpaRTA (0.048%) 0.12 95.6 0.31 88.3 0.13 95.6 0.23 90.1 0.25 90.9 0.30 88.6 0.35 65.2 0.26 89.7 88.0 SpaRTA (0.048%) O,V 0.11 95.9 0.31 88.4 0.14 95.8 0.23 90.4 0.24 91.2 0.30 88.4 0.34 64.9 0.26 89.7 88.1
+Table 6: Test loss and accuracy of the mistral-7b-it model adapted to GLUE and BoolQ datasets with different fine-tuning methods. For training details see Appendix E.2.
+MMLU gemma-2b gemma-2b-it mistral-7b mistral-7b-it loss acc.
+loss acc.
+loss acc.
+loss acc.
+PT (zero-shot) - - 5.284 35.3% - - 1.839 59.3% SpaRTA 1.249 45.1% 1.250 45.1% 0.987 61.5% 0.930 63.1% LoRA 1.233 46.9% 1.271 45.1% 0.981 62.8% 0.928 63.0%
+Table 7: Test loss and accuracy of models adapted to the MMLU dataset with SpaRTA and LoRA.
+We also report zero-shot performance of instruction following PT models.
+For training details see Table 17 in the Appendix.
+We now focus on comparing SpaRTA and LoRA methods using the 7 NLU task in GLUE tasks as well as BoolQ.
+Table 5 presents the results for adapting gemma-2b-it.
 An equivalent set of results for the gemma-2b base model is given in Table 10, see Appendix D.
-Here again, adaptation results (rows) are ordered in descending order of density with SpaRTA evaluated for different orders of magnitude of density levels 5%, 0.5%, 0.037% – the last one allowing SpaRTA to have approximately the same number of trainable parameters as LoRA.
-LoRA has rank $r=8$ equivalent to 0.037% of trainable parameters; Head adaptation uses a very small 1.63e-4% density.
-Overall, the same trend as for IMDB is observed.
-Full FT results are very often the best, sometimes bested by SpaRTA 5% (RTE, MNLI).
-For a low density, where LoRA and SpaRTA use the same number of trainable parameters, SpaRTA can best LoRA (QNLI, SST2, RTE) but is overall competitive with LoRA results.
-A similar set of results on the same GLUE tasks is provided for mistral-7b-it in Table 5.
-Again, results for the mistral-7b PT base model can be found in Table 11, see Appendix D.
-Full FT results can often be matched or bested by SpaRTA 5% (RTE, SST2, QQP, MNLI, MRPC, COLA).
-SpaRTA often gives better results than LoRA (QNLI, RTE, SST2, QQP, MNLI) for a low density (0.0048%) where both techniques have comparable numbers of parameters.
-Once again, SpaRTA is a competitive adaptation technique, often providing similar results to LoRA with low density.
+Adaptation results (rows) are grouped by which type of parameters are targeted, with SpaRTA further ordered in descending order of density.
+SpaRTA with 0.037% density has approximately the same number of trainable parameters as LoRA, for which a rank $r=8$ was used.
+When targeting all parameters, SpaRTA overall exhibits better performance with higher density, i.e, when more parameters are chosen to be trained, with a few exceptions.
+Targeting the Wq and Wv self-attention matrices shows very similar performance between LoRA, DoRA, and SpaRTA.
+Targeting the Wo and Wv self-attention matrices within SpaRTA, which was shown to be the optimal in Section 7.2, yields better performance for SpaRTA when compared to both LoRA and DoRA on four of the eight datasets (QNLI, RTE, MRPC, BoolQ).
+A similar set of results on GLUE and BoolQ is provided for mistral-7b-it in Table 6.
+Again, results for the mistral-7b base model can be found in Table 11 in Appendix D.
+Again, SpaRTA exhibits comparable performance to LoRA and DoRA for a low density of 0.0048%, where all methods have comparable numbers of trainable parameters.
+And once again, with SpaRTA targeting Wv and Wo typically outperforming the targeting of Wq and Wv.
+7.4 MMLU
+We have chosen MMLU because it is known for being a very challenging NLU task.
+Table 7 compares SpaRTA against LoRA on adapting each of our models to the MMLU multiple choice question answering task, where each model must predict the correct answer to a set of questions.
+Solving this task requires a high level of world knowledge and problem solving skills.
+In this experiment, we restrict both methods to use approximately the same number of trainable parameters.
+See Table 17 in the Appendix for training details.
+Here again, the results show that SpaRTA is competitive with LoRA, providing similar performance.
 7.5 Remarks
-Results on GLUE and IMDB establish that SpaRTA is a viable adaptation technique that can be competitive with LoRA, especially for larger LMs.
+Our results establish that SpaRTA is a viable adaptation technique that can be competitive with LoRA, especially for larger LMs.
 These results demonstrate that a simple sparsifying scheme can offer a valid adaptation technique.
-This opens up the possibility for more investigation in sparse adaptation as the low rank approximation of a model’s parameter changes is not the only mechanism to provide a performant adaptation method.
-This indicates that what matters is not necessarily the adapter structure used in PEFT techniques, but rather the number of independent parameters, that is relevant to the adaptation task at hand.
-8 Exploring Model Merging
-Yu et al.
-(2024) demonstrate that one can perform Full FT, compute $\Delta=\theta_{\text{FT}}-\theta_{\text{PT}}$, sparsify $\Delta$ to get $\tilde{\Delta}$, and obtain good performance using $\tilde{\theta}_{\text{FT}}=\theta_{\text{PT}}+\tilde{\Delta}$ as model parameters.
-Furthermore, they show that merging models via their sparsified $\tilde{\Delta}$-s from different tasks can also maintain the performance of the individual models.
-This motivates exploring model merging via SpaRTA, where no Full FT is required and sparse adapters are instead fine-tuned.
-In our setup, because the heads of our SpaRTA-tuned models are fully trained, merging them into a single head would create parameter interference.
-We thus merge everything except the heads and add each unmerged head on top of the merged model creating a multi-head model.
-We need the merged model to produce multiple outputs (i.e., one for each task) per input and this is exactly what a multi-head merged model does.
-In the forward pass, a sequence of input tokens goes through the merged model, producing a final state representation of the tokens.
-This representation is shared by all tasks and is passed to each head to produce one prediction per task as output.
-This way all tasks are processed concurrently on the same input.
-We experiment merging two models: one individually fine-tuned with SpaRTA for SST2 (classifying text sentiment) and another for COLA (classifying grammatical correctness), using gemma-2b as PT model and 99% sparsity in both cases.
-We merge these two models by adding the respective $\Delta_{\phi}$ to their PT model’s original parameters $\theta_{\text{PT}}$, and obtaining a two-headed model that can simultaneously classify any input sentence on both criteria in a single forward pass.
-Table 6 compares performance of the merged model with that of the unmerged models.
-This result encourages future exploration of this practical usecase for SpaRTA.
-SST2 COLA
-Model loss acc loss acc
-SST2 0.113 96.3% - -
-COLA - - 0.487 82.7%
-Merged 0.118 96.7% 0.428 84.4%
-Table 6: Test loss an accuracy of SST2 and COLA models fine-tuned on gemma-2b using SpaRTA with 99% sparsity. The merged model outperforms both individual models in this scenario.
+This opens the possibility for further investigating sparse adaptation as the low rank approximation of a model’s parameter changes is not the only mechanism to provide a performant adaptation method.
+This indicates that what matters is not necessarily the adapter structure used in PEFT techniques, but rather the number of trainable parameters, that is relevant to the adaptation task at hand.
 8 Conclusion
@@
 Experiments with 2B and 7B parameter pre-trained models demonstrate good performance, and as per Yu et al.
-(2024), we expect larger models to allow for higher levels of sparsity in training; meaning that efficiency of SpaRTA should get better with larger model sizes as also suggested in Table 1.
-Regarding future directions, while SpaRTA has been applied to supervised learning, it is also amenable to reinforcement learning often used for model alignment (Ouyang et al., 2022a).
-Furthermore, SpaRTA allows one to select which modules to sparsify and at what layers.
-Like LoRA, this decision affects both the number of trainable parameters and the performance of the fine-tuned model, and we plan to investigate its impact in SpaRTA.
-Lastly, we plan to explore model merging as discussed in Section 8; SpaRTA opens the potential for model merging without Full FT but with possibly little interference.
+(2024), we expect larger models to allow for higher levels of sparsity in training, meaning that efficiency of SpaRTA should get better with larger model sizes as also suggested in Table 1.
+Regarding future directions, while SpaRTA has been applied to supervised learning, it is also amenable to reinforcement learning often used for model alignment (Ouyang et al., 2022).
+We also plan to explore the merging of SpaRTA adapters due to their potential for little interference.
 9 Limitations
@@
 We compute all delta weight matrices for two well-known instruction-following fine-tuned models: gemma-2b-it and mistral-7b-it, see Section 6.2.
-As we can see in Tables 7 and 8, the feedforward (MLP) delta $\Delta$-matrices associated with each layer of the transformer network are all full rank in both models.
+As we can see in Tables 8 and 9, the feedforward (MLP) delta $\Delta$-matrices associated with each layer of the transformer network are all full rank [^5] We compute the rank of a matrix as the number of singular values that are greater than zero, given a specified tolerance, using torch.linalg.matrix_rank.
+We did some sensitivity analysis to determine such tolerance.
+in both models.
 This is also the case for all self attention key (k) and value (v) projection matrices.
-The self attention query (q) and output (o) $\Delta$-matrices all show relatively small rank deficiencies: between $9$ and $2$ for the query projection matrices and between $46$ and $3$ for the output projection matrices, out of a potential maximum rank of $2,048$ for the gemma-2b-it model; and between $1,788$ and $28$ for the query projection matrices and between $246$ and $10$ for the output projection matrices, out of a potential maximum rank of $4,096$ for mistral-7b-it model.
-The token embedding $\Delta$-matrix for mistral-7b-it is full rank.
-However, it has a rank of $97$ for gemma-2b-it, being the only matrix for that model whose delta change shows a significant rank deficiency (of $1,951$).
-Of the two models, mistral-7b-it is the only one that does not tie its tokens embedding weights to its head; so that mistral-7b-it model’s head (untied LM head), which is trained independently from the tokens embedding, has a rank deficiency of $1$ for its delta change.
-Basically, we observe that the fine-tuning changes in all weight matrices of these models are full rank or very close to it, with the only exception being the changes in the tokens embedding for gemma-2b-it which are significantly low rank.
-We analyzed here the changes in fine-tuned weight matrices for two popular open-weights language models.
-As discussed in Section 2.1, we compared instruction-following models with their original, pre-trained versions, by computing the differences (delta) in their weight matrices.
-Tables 7 and 8 provide a detailed breakdown of the ranks for all these delta matrices.
+However, the self attention query (q) and output (o) $\Delta$-matrices all show relatively small rank deficiencies: between $9$ and $2$ for the query projection matrices and between $46$ and $3$ for the output projection matrices, out of a potential maximum rank of $2,048$ for the gemma-2b-it model; and between $1,788$ and $28$ for the query projection matrices and between $246$ and $10$ for the output projection matrices, out of a potential maximum rank of $4,096$ for the mistral-7b-it model.
+The token embedding $\Delta$-matrices of both the gemma-2b-it and mistral-7b-it models are full rank.
+Of the two models, mistral-7b-it is the only one that does not tie its token embeddings’ weights to its head; thus the mistral-7b-it’s model head (untied lm head) is updated independently during fine-tuning, and the resulting delta change of this matrix is also full rank.
+Basically, we observe that the fine-tuning changes in the weight matrices of these well known models are all full rank, with the only exception being the changes in the query (q) and output (o) projection matrices that show small rank deficiencies.
 The observed ranks suggest that constraining the (delta) changes in weight matrices to be low-rank is not essential for fine-tuning models efficiently.
-Weight matrix dims Layer(s) Rank Rank deficiency
-tokens embed [ 256000, 2048 ] - 97 1951
-self attn q proj [ 2048, 2048 ] 0 2039 9
-15 2041 7
-14 2042 6
-6 2044 4
-1,3-5,7,8,10-13,16,17 2045 3
-2,9 2046 2
+Weight matrix dims Layer(s) Rank(s) Rank deficiencies
+tokens embed [ 256000, 2048 ] - 2048 0 (full rank)
+self attn q proj [ 2048, 2048 ] 0 - 17 2046 - 2039 2 - 9
 self attn k proj [ 256, 2048 ] 0 - 17 256 0 (full rank)
 self attn v proj [ 256, 2048 ] 0 - 17 256 0 (full rank)
-self attn o proj [ 2048, 2048 ] 0 2002 46
-16 2040 8
-11,12,14 2042 6
-6,8,10,13,15 2043 5
-3-5,7,9,17 2044 4
-1,2 2045 3
+self attn o proj [ 2048, 2048 ] 0 - 17 2045 - 2002 3 - 46
 mlp gate proj [ 16384, 2048 ] 0 - 17 2048 0 (full rank)
@@
 Table 8: Rank of the change differences in gemma-2b-it model weight matrices after full fine-tuning.
-Weight matrix dims Layer(s) Rank Rank deficiency
+Weight matrix dims Layer(s) Rank(s) Rank deficiencies
 tokens embed [ 32768, 4096 ] - 4096 0 (full rank)
-self attn q proj [ 4096, 4096 ] 0 2308 1788
-1 3103 993
-3 4023 73
-31 4046 50
-30 4050 46
-2,4 4051 45
-6 4052 44
-29 4056 40
-28 4057 39
-11,27 4058 38
-8,9 4059 37
-10,19,25,26 4061 35
-12,18,23 4062 34
-5,7,14,17 4063 33
-13,20,22 4064 32
-15 4065 31
-21,24 4066 30
-16 4068 28
+self attn q proj [ 4096, 4096 ] 0 - 31 4068 - 2308 28 - 1788
 self attn k proj [ 1024, 4096 ] 0 - 31 1024 0 (full rank)
 self attn v proj [ 1024, 4096 ] 0 - 31 1024 0 (full rank)
-self attn o proj [ 4096, 4096 ] 0 3850 246
-31 4059 37
-1,30 4060 36
-14 4068 28
-18 4069 27
-23 4070 26
-2,10,26 4071 25
-8,27 4072 24
-3,9 4074 22
-4,11,29 4075 21
-6,13,22,28 4076 20
-7,12 4077 19
-15,16,19 4079 17
-17 4080 16
-5 4081 15
-25 4083 13
-21,24 4085 11
-20 4086 10
+self attn o proj [ 4096, 4096 ] 0 - 31 4086 - 3850 10 - 246
 mlp gate proj [ 14336, 4096 ] 0 - 31 4096 0 (full rank)
@@
 mlp down proj [ 4096, 14336 ] 0 - 31 4096 0 (full rank)
-untied lm head [ 4096, 32768 ] - 4095 1
+untied lm head [ 4096, 32768 ] - 4096 0 (full rank)
 Table 9: Rank of the change differences in mistral-7b-it model weight matrices after full fine-tuning.
 Appendix B Additional Dataset details
-Dataset Classes Train Dev Test
-COLA 2 7,551 1,000 1,043
-MNLI 3 100,000 10,000 19,647
-MRPC 2 3,668 408 1,725
-QNLI 2 99,701 5,000 5,463
-QQP 2 100,000 5,000 40,430
-RTE 2 2,165 300 277
-SST-2 2 66,349 1,000 872
-IMDB 2 25,000 5,000 20,000
-Table 9: Sequence classification datasets. Training sets limited to 100K samples. Training samples with more than 256 tokens are removed (here using mistral-7b tokenizer).
-IMDB: contains a sample of “highly polar” movie reviews obtained from the online Internet Movie Database (IMDb) website.
+IMDB contains a sample of “highly polar” movie reviews obtained from the online Internet Movie Database (IMDb) website.
 IMDb registered users provide a rating (from $1$ to $10$) with each review.
@@
 No reviews with ratings beyond these ranges are present in the dataset, which was constructed to have an equal number of positive and negative reviews, so guessing randomly yields 50% accuracy.
-GLUE: Refer to Wang et al. (2019) which contains descriptions and details on all GLUE datasets.
-Table 9 shows the actual splits used in our experiments for the IMDB and GLUE datasets, where the training data was filtered using the mistral-7b tokenizer by a threshold in the maximum token length of an network input.
-The only differences are for the training splits of the MRPC, QNLI, and RTE datasets.
+GLUE datasets are described in details in Wang et al. (2019).
+BoolQ Clark et al.
+(2019) is a reading comprehension dataset with binary yes/no questions.
+Each example is a triplet of (passage, question, answer).
+It represents a natural language inference task in which a passage and a question are given as input, and a yes or no answer should be predicted.
+MMLU Hendrycks et al.
+(2020) is a benchmark designed to measure the knowledge a model acquires during pre-training.
+To facilitate the elicitation of a model’s knowledge using multiple-choice questions, MMLU also comes with a dataset of examples on multiple-choice question answering.
+Thus, this training set is not designed to increase the model knowledge about the world, but to teach a model to answer knowledge questions in multiple-choice format in a zero-shot setting.
 Appendix C Training details
@@
 We only do this for the training data since evaluation (on the development or test data) requires much less compute and memory (no gradients need to be calculated and stored) and is performed less frequently.
-We will indicate in Tables 2 and 9 the final splits after filtering the training data this way, with the final number of training examples used for the fine-tuning.
+We indicate in Tables 2 the final splits after filtering the training data this way, with the final number of training examples used for the fine-tuning.
 With SpaRTA, we froze the token embeddings layer, made fully trainable the classification head, and randomly chose a sparse proportion of (scalar) parameters to be trained in all the other layers of a model.
@@
 Appendix D Additional Experiment Results
-Tables 10 and Table 11 present the complete set of results on GLUE for the PT base models gemma-2b and mistral-7b, respectively.
-Model: google/gemma-2-2b QNLI RTE SST2 QQP MNLI MRPC COLA loss acc.
+Tables 10 and Table 11 present the complete set of results on GLUE and BoolQ for the PT base models gemma-2b and mistral-7b, respectively.
+Tables 12 and 13 present standard errors from corresponding experiments.
+Model: gemma-2b QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ Method Targets loss acc.
 loss acc.
@@
 loss acc.
-loss mcc Full FT 0.20 92.6 0.43 80.1 0.11 96.7 0.28 88.3 0.36 86.4 0.33 86.1 0.41 60.3 SpaRTA (5%) 0.15 94.4 0.46 80.5 0.11 96.8 0.23 90.1 0.28 89.4 0.38 83.4 0.34 64.6 SpaRTA (0.5%) 0.15 94.1 0.61 69.0 0.12 96.3 0.24 89.8 0.30 89.0 0.38 83.4 0.38 62.2 SpaRTA (0.037%) 0.29 88.3 0.69 57.8 0.11 96.2 0.30 86.9 0.47 81.1 0.58 70.3 0.52 42.1 LoRA (0.037%) 0.17 93.6 0.59 69.0 0.11 96.1 0.26 88.6 0.30 88.6 0.36 84.5 0.36 61.7 Head (1.63e-4%) 0.56 71.8 0.68 60.3 0.32 89.1 0.43 79.9 0.89 58.7 0.59 70.4 0.53 33.2
-Table 10: Test loss and accuracy of model google/gemma-2-2b adapted to glue datasets with different fine-tuning methods. For training details see Table 13.
-Model: mistralai/Mistral-7B-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA loss acc.
+loss mcc loss acc.
+SpaRTA (5%) ALL 0.15 94.5 0.50 80.5 0.12 96.6 0.23 90.2 0.39 85.5 0.38 84.6 0.36 62.6 0.32 86.7 SpaRTA (0.5%) 0.15 94.5 0.51 77.3 0.12 96.1 0.23 90.1 0.42 84.1 0.37 84.3 0.36 62.3 0.34 85.4 SpaRTA (0.037%) 0.20 91.9 0.62 66.3 0.13 95.4 0.27 88.3 1.34 49.3 0.51 76.1 0.49 50.7 0.39 82.8 SpaRTA (0.037%) Q,V 0.18 93.0 0.58 70.4 0.13 95.7 0.26 88.8 1.04 56.8 0.43 80.3 0.41 56.9 0.36 84.5 LoRA (0.037%) 0.15 94.2 0.62 65.5 0.11 96.8 0.26 88.9 0.42 84.0 0.47 77.3 0.40 60.9 0.34 85.7 DoRA (0.037%) 0.15 94.4 0.62 66.5 0.11 96.6 0.25 89.0 0.41 84.4 0.47 77.6 0.39 60.6 0.43 78.9 SpaRTA (0.037%) O,V 0.16 93.8 0.55 73.2 0.11 96.5 0.25 89.3 0.51 80.4 0.39 84.3 0.39 61.9 0.34 86.1
+Table 10: Test loss and accuracy of model gemma-2b adapted to GLUE and BoolQ datasets with different fine-tuning methods.
+Results are averaged over 3 random seeds.
+For training details see Appendix E.2.
+Model: Mistral-7B-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method Targets loss acc.
 loss acc.
@@
 loss acc.
-loss mcc Full FT 0.12 95.4 0.28 89.9 0.11 96.9 0.23 90.1 0.25 91.1 0.34 85.9 0.32 66.8 SpaRTA (5%) 0.12 95.5 0.34 86.3 0.11 96.9 0.23 90.3 0.25 90.9 0.34 87.6 0.37 64.3 SpaRTA (0.5%) 0.12 95.9 0.35 86.3 0.11 96.7 0.22 90.5 0.24 91.2 0.32 88.8 0.34 68.2 SpaRTA (0.048%) 0.13 95.3 0.36 86.3 0.11 96.8 0.24 89.7 0.26 90.3 0.38 86.3 0.34 66.9 LoRA (0.048%) 0.12 95.6 0.30 88.1 0.11 96.7 0.23 90.0 0.25 91.0 0.31 87.3 0.32 67.8 Head (1.15e-4%) 0.39 82.8 0.63 70.4 0.16 94.8 0.33 85.8 0.63 74.0 0.53 73.8 0.48 45.5
-Table 11: Test loss and accuracy of model mistralai/Mistral-7B-v0.3 adapted to glue datasets with different fine-tuning methods. For training details see Table 13.
+loss mcc loss acc.
+SpaRTA (5%) ALL 0.11 95.9 0.36 85.0 0.11 96.8 0.38 82.1 0.30 89.2 0.33 86.7 0.35 64.6 0.25 90.1 SpaRTA (0.5%) 0.12 95.8 0.44 84.4 0.11 96.7 0.22 90.6 0.26 90.4 0.32 87.0 0.35 66.8 0.26 89.9 SpaRTA (0.048%) 0.13 95.3 0.41 83.6 0.11 96.7 0.23 90.1 0.40 86.7 0.34 86.6 0.36 64.6 0.26 89.8 SpaRTA (0.048%) Q,V 0.14 94.5 0.37 84.0 0.11 96.6 0.24 89.7 0.33 88.6 0.36 85.5 0.36 65.2 0.28 89.3 LoRA (0.048%) 0.12 95.4 0.39 85.4 0.12 96.5 0.23 90.1 0.28 89.7 0.32 87.0 0.38 64.5 0.27 89.6 DoRA (0.048%) 0.12 95.6 0.41 85.3 0.12 96.8 0.24 90.1 0.28 89.7 0.33 86.4 0.34 65.8 0.26 89.9 SpaRTA (0.048%) O,V 0.15 94.5 0.41 85.6 0.21 91.7 0.24 89.6 0.33 87.8 0.35 86.4 0.39 59.8 0.29 89.3
+Table 11: Test loss and accuracy of model mistral-7b adapted to GLUE and BoolQ datasets with different fine-tuning methods.
+Results are averaged over 3 random seeds.
+For training details see Appendix E.2.
+Model: gemma-2b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc.
+loss acc.
+loss acc.
+loss acc.
+loss acc.
+loss acc.
+loss mcc loss acc.
+PT (zero-shot) 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 SpaRTA (5%) ALL 0.00 0.14 0.03 1.54 0.01 0.51 0.00 0.21 0.01 0.26 0.01 0.48 0.00 0.88 0.01 0.54 SpaRTA (0.5%) 0.00 0.16 0.03 0.64 0.01 0.46 0.00 0.17 0.00 0.13 0.01 0.07 0.01 0.75 0.01 0.25 SpaRTA (0.037%) 0.00 0.12 0.01 0.55 0.00 0.17 0.00 0.18 0.00 0.14 0.00 0.45 0.01 0.44 0.00 0.09 LoRA (0.037%) Q,V 0.00 0.11 0.01 1.10 0.00 0.35 0.00 0.03 0.00 0.11 0.00 0.22 0.02 1.40 0.01 0.12 DoRA (0.037%) 0.00 0.11 0.01 0.55 0.00 0.20 0.00 0.03 0.00 0.09 0.00 0.32 0.01 1.31 0.01 0.29 SpaRTA (0.037%) 0.00 0.08 0.00 1.04 0.01 0.11 0.00 0.00 0.00 0.26 0.00 0.22 0.02 2.83 0.00 0.41 SpaRTA (0.037%) O,V 0.00 0.13 0.01 0.64 0.00 0.27 0.00 0.16 0.00 0.17 0.02 0.22 0.01 4.36 0.00 0.22
+Table 12: Standard errors of test loss and accuracy of model gemma-2b-it adapted to GLUE and BoolQ datasets with different fine-tuning methods.
+Model: Mistral-7B-Instruct-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc.
+loss acc.
+loss acc.
+loss acc.
+loss acc.
+loss acc.
+loss mcc loss acc.
+PT (zero-shot) 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 SpaRTA (5%) ALL 0.00 0.09 0.01 0.24 0.01 0.26 0.00 0.03 0.00 0.08 0.01 0.47 0.01 1.91 0.00 0.16 SpaRTA (0.5%) 0.00 0.05 0.01 1.07 0.01 0.18 0.00 0.10 0.00 0.09 0.02 0.25 0.01 1.28 0.00 0.09 SpaRTA (0.048%) 0.00 0.10 0.03 0.55 0.00 0.20 0.00 0.03 0.00 0.11 0.01 0.32 0.02 1.87 0.00 0.20 LoRA (0.048%) Q,V 0.00 0.16 0.01 0.60 0.01 0.44 0.00 0.06 0.00 0.13 0.01 0.33 0.01 1.16 0.00 0.17 DoRA (0.048%) 0.02 0.76 0.02 0.87 0.00 0.17 0.00 0.03 0.00 0.07 0.00 0.17 0.01 1.22 0.00 0.16 SpaRTA (0.048%) 0.00 0.06 0.01 0.73 0.00 0.23 0.00 0.04 0.00 0.05 0.01 0.43 0.01 1.17 0.00 0.27 SpaRTA (0.048%) O,V 0.00 0.02 0.01 0.42 0.01 0.34 0.00 0.07 0.00 0.13 0.02 0.24 0.01 0.97 0.00 0.23
+Table 13: Standard errors of test loss and accuracy of model mistral-7b-it adapted to BLUE and BoolQ datasets with different fine-tuning methods.
 Appendix E Training Hyper-Parameters
@@
 That is 77%, 75%, 73%, and 71% of the original training data of 25,000 examples, respectively.
+Model: google/gemma-2b-it method QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ SpaRTA (5%) 1e-5 5e-4 5e-5 5e-5 1e-5 2e-4 5e-5 5e-5 SpaRTA (0.5%) 2e-4 1e-3 2e-4 2e-4 2e-4 5e-4 1e-3 2e-4 SpaRTA (0.037%) 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 LoRA (0.037%) 2e-4 1e-3 5e-4 5e-4 5e-4 2e-4 5e-4 5e-4 DoRA (0.037%) 2e-4 1e-3 5e-4 5e-4 5e-4 2e-4 5e-4 5e-4 SpaRTA (0.037%) 5e-4 1e-3 5e-4 5e-4 5e-4 5e-4 1e-3 5e-4 SpaRTA (0.037%) 5e-4 1e-3 5e-4 5e-4 5e-4 5e-4 1e-3 5e-4
+Table 14: Optimal learning rates for gemma-2b-it model for different fine-tuning methods.
+Model: Mistral-7B-Instruct-v0.3 method QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ SpaRTA (5%) 1e-5 5e-6 5e-6 1e-5 1e-5 5e-5 5e-5 5e-6 SpaRTA (0.5%) 2e-4 1e-4 5e-5 5e-4 2e-4 5e-4 2e-4 5e-5 SpaRTA (0.048%) 5e-4 5e-4 5e-4 1e-3 5e-4 1e-3 1e-3 1e-3 LoRA (0.048%) 2e-4 1e-4 2e-4 2e-4 1e-4 2e-4 2e-4 1e-4 DoRA (0.048%) 2e-4 1e-4 2e-4 2e-4 1e-4 2e-4 2e-4 1e-4 SpaRTA (0.048%) 2e-4 1e-4 2e-4 5e-4 2e-4 1e-3 1e-3 5e-4 SpaRTA (0.048%) 2e-4 1e-4 2e-4 5e-4 2e-4 1e-3 1e-3 5e-4
+Table 15: Optimal learning rates for mistral-7b-it model for different fine-tuning methods.
 Parameter gemma-2b gemma-2b-it mistral-7b mistral-7b-it Full FT batch size 32 32 36 36 num epochs 2 2 $\ast$ $\ast$ learning rate 1e-5 1e-5 3e-6 3e-6 max grad norm 10 50 120 120 dropout 0.1 0.1 0.15 0.15 weight decay 0.1 0.1 0.01 0.01 SpaRTA batch size 40 40 16 16 num epochs.
@@
 For SpaRTA, parameters for density 5%, 0.5% 0,05% are reported.
-E.2 GLUE
-For GLUE, given the number of tasks (7) and the number of methods tried (6), reporting each best set of training parameters would be just prohibitive.
-Instead, the trainings of these methods was done using a simple grid search over some of the hyper-parameters.
-The range on these hyper-parameters are reported in Table 13.
+E.2 GLUE and BoolQ
+For GLUE and BoolQ, training was done using a simple grid search over the learning rates: [1e-3, 5e-4, 2e-4, 1e-4, 5e-5, 1e-5, 5e-6].
+The optimal learning rates found are given in Tables 14 and 15.
+Weight decay was set to 0.
+Dropout was set to 0.0 for SpaRTA and 0.1 for LoRA and DoRA.
+Batch sizes were set according to what could fit in GPU memory (different per dataset and model).
+The number of epochs was set to 3 for instruct models and 2 for base models.
 Similarly to IMDB, any sample longer than 256 tokens (which is dependnent on the tokenizer used) is discarded from the dataset to avoid having few minibatch with one very long sample compared to the rest, as seen in Table 2.
-Parameter Grid Search Value Range
-Full FT batch size [32, 24, 8] depending on task and model to fit in GPU memory
-learning rate [1e-3, 5e-4, 2e-4, 1e-4, 5e-5, 1e-5, 5e-6]
-dropout [0.1]
-weight decay [0.1]
-SpaRTA batch size [same as for Full FT]
-learning rate [same as for Full FT]
-weight decay [0.1]
-LoRA batch size [same as for Full FT]
-learning rate [same as for Full FT]
-$r$ [8]
-$\alpha$ [16]
-dropout [0.1]
-weight decay [0.1]
-Head batch size [same as for Full FT]
-learning rate [same as for Full FT]
-dropout [0.1]
-weight decay [0.0]
-Table 13: (GLUE) Training parameters and their grid search value range to get results for each adaptation method for each model in Table 4, Table 5, Table 10, and Table 11.
-Best set of parameters were chosen from best validation loss results over 2 epochs of training.
+E.3 MMLU
+As with other datasets, we started approximating the deltas of the self-attention matrices Wq and Wv with rank $r=8$ matrices, resulting in approximately 1M trainable parameters for the Gemma 2B models and 3.5M for the Mistral 7B models under LoRA ($\alpha=16$).
+We choose the sparsity of our SpaRTA adapter accordingly so both methods end up with the same number of parameters to train.
+Thus, our SpaRTA approach uses a $s=99.96\%$ for the Gemma models (i.e.
+making only $0.04\%$ of the original parameters trainable); and $s=99.95\%$ for the Mistral models.
+Since we observed that gemma-2b struggled to learn with both SpaRTA and LoRA methods, we decided to increase the number of trainable parameters when adapting gemma-2b to MMLU.
+Specifically, we increased the rank of the LoRA adaption matrices to $r=16$, which lead to approximately 2M trainable parameters; and chose accordingly a sparsity $s=99.92\%$ for SpaRTA.
+The MMLU dataset has a small fraction of training examples that are extremely long.
+We enforce a maximum input token length of $520$ for training efficiency.
+This reduces the number of training examples from $99,842$ to $84,296$ and $74,100$ for the Gemma and Mistral instruction models; and to $91,321$ and $85,820$ for their respective base models.
+Test and validations sets are not affected by this decision.
+The training parameters used for adapting each model to MMLU are shown in Table 17.
+Training parameter gemma-2b gemma-2b-it mistral-7b mistral-7b-it
+SpaRTA  / LoRA epochs 14 / 6 1 1 1
+batch size 40 40 40 40
+learning rate 1e-4 1e-4 5e-5 5e-5
+dropout 0.0 0.05 0.1 0.1
+weight decay 0.0 0.0 0.0 0.0
+Table 17: Training parameters used with both SpaRTA and LoRA for each of the pre-trained models in Table 7 (MMLU).

### v3

@@
 Parameter-Efficient Fine-Tuning (PEFT) methods aim at significantly reducing the computational and memory resources needed for fine-tuning these models by only training on a small number of parameters instead of all model parameters.
-Currently, the most popular PEFT method is the Low-Rank Adaptation (LoRA), which freezes the parameters of the model to be fine-tuned and introduces a small set of trainable parameters in the form of low-rank matrices.
-We propose simply reducing the number of trainable parameters by randomly selecting a small proportion of the model parameters to train on.
-In this paper, we compare the efficiency and performance of our proposed approach with PEFT methods, including LoRA, as well as full parameter fine-tuning.
+Currently, the most popular PEFT method is the Low-Rank Adaptation (LoRA), which freezes the parameters of the model and introduces a small set of trainable parameters in the form of low-rank matrices.
+We propose simply reducing the number of trainable parameters by randomly selecting a small proportion of the model parameters to train on, while fixing all other parameters, without any additional prior assumptions such as low-rank structures.
+In this paper, we compare the efficiency and performance of our proposed approach to other PEFT methods as well as full parameter fine-tuning.
+We find our method to be competitive with LoRA when using a similar number of trainable parameters.
+Our findings suggest that what truly matters for a PEFT technique to perform well is not necessarily the specific adapter structure, but rather the number of trainable parameters being used.
 Sparsity May Be All You Need:
@@
 The idea behind PEFT (Lialin et al., 2023) is to significantly reduce the number of trainable parameters during fine-tuning while maintaining performance.
-Figure 1: The proposed SpaRTA method which randomly chooses a subset of parameter indices vector $\phi$ from pre-trained model parameters $\theta_{\text{PT}}$ and updates them via adapter $\Delta_{\phi}$.
+Figure 1: The proposed SpaRTA method which randomly chooses a small subset of parameters from $\theta_{\text{PT}}$, stores the indices of the selected parameters in $\phi$ and updates the model via adapter $\Delta_{\phi}$.
 Low Rank Adaptation (LoRA), first introduced by Hu et al.
 (2022), currently remains the most popular PEFT technique.
-LoRA freezes all the pre-trained model parameters $\theta_{\text{PT}}$ and introduces trainable low-rank matrices (e.g., $B$, $A$) for a pre-selected subset of matrix tensors to represent the changes ($\Delta=BA$) needed for adapting the model to a new task.
+LoRA freezes all the pre-trained model parameters $\theta_{\text{PT}}$ and introduces trainable low-rank matrices (e.g., $B$, AA) to represent the changes ($\Delta=BA$) needed for adapting the model to a new task.
 The adapted model parameters are given by $\theta_{\text{PT}}+\Delta$.
@@
 See Figure 1 for an illustration of the method.
-Generally, one samples $m$ parameters from the pre-trained model $\theta_{\text{PT}}$, stores the indices of these parameters in $\phi$, and uses adapter $\Delta_{\phi}$ to fine-tune $\theta_{\text{PT}}$.
+Generally, one samples $m$ parameters from the pre-trained model $\theta_{\text{PT}}$, stores their indices in $\phi$, and uses adapter $\Delta_{\phi}$ to fine-tune $\theta_{\text{PT}}$.
 To investigate the performance of SpaRTA, we build adapters with different sparsity levels and evaluate them on a wide range of natural language understanding benchmarks.
@@
 (2024) look at the differences between a language model’s parameters before and after fine-tuning, and demonstrate empirically that it is possible to randomly drop (i.e., set to zero) up to 99% of these parameter changes, represented by $\Delta$, without significantly affecting model performance.
-This motivates our approach, SpaRTA, which produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters: SpaRTA randomly selects the (scalar) parameters to train and freezes the remaining parameters (i.e., setting the corresponding $\Delta$ values to zero).
+This motivates our approach, SpaRTA, which produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters.
+SpaRTA randomly selects the (scalar) parameters to train and freezes the remaining parameters (i.e., setting the corresponding $\Delta$ values to zero).
 This $\Delta$-sparsity also helps in reducing overfitting, as pre-trained models typically have more than enough capacity to learn the often limited amount of (labeled) data used for fine-tuning.
@@
 SpaRTA, like LoRA, reduces the number of gradient computations during training (when compared with full fine-tuning), and ultimately has the same inference cost as the original model (after merging back the sparse $\Delta$ into the model).
-The properties of low memory, low computation needs in training, identical inference time as the original model are all quite desirable for an adaptation technique.
+For an adaptation technique, having lower memory and computation needs during training as well as identical inference time to the original model are all quite desirable properties.
 SpaRTA has all these properties plus the unique added benefit of producing only sparse changes in a small number of the parameters of the original model that can be beneficial for merging multiple SpaRTA adapters.
 2.1 Is Low-Rank Adaption Necessary?
-In Appendix A, we show that the changes (i.e., $\Delta$) in weight matrix during full parameter fine-tuning are, in fact, not generally low-rank for capable models such as gemma-2b and mistral-7b.
-This indicates that LoRA works, not particularly because of its low-rank constraint, but rather due to the capacity reduction achieved since LoRA fine-tuning with is typically done on limited training data.
-Such insight also hints that any constraint which reduces the capacity of the original model could perform competitively, motivating our approach that selects a small number of parameters of the original model to be updated during training.
+In Appendix A, we show that the changes (i.e., $\Delta$) in weight matrix during full parameter fine-tuning are, in fact, not generally low-rank for capable models such as gemma-2b-it and mistral-7b-it.
+This indicates that LoRA works, not particularly because of its low-rank constraint, but rather due to the reduction in model capacity achieved by LoRA when fine-tuning on limited training data, as it is typically done in task adaption.
+Such insight also hints that any constraint reducing the capacity of the original model could perform competitively, motivating our approach which selects a small number of parameters from the original model to be updated during training.
 3 Related Work
-The last few years has seen many advances in fine-tuning Large Language Models (LLMs).
+The last few years has seen many advances in PEFT methods.
 Perhaps the most well-known and used method in practice is LoRA Hu et al.
@@
 4 SpaRTA: Adapting a Random Subset of Model Parameters
-Suppose the vectorized parameters of the pre-trained language model are $\theta_{\text{PT}}\in\mathbb{R}^{n}$ where $n$ is the number of parameters, and full parameter fine-tuning (FT) of the model is performed with a labeled dataset characterizing a task.
+Suppose the parameters of a pre-trained language model are $\theta_{\text{PT}}\in\mathbb{R}^{n}$ where $n$ is the number of parameters, and full parameter fine-tuning (FT) is performed with a labeled dataset characterizing a task.
 FT typically updates $\theta_{\text{PT}}$ using a stochastic first-order gradient-based optimization algorithm, e.g., Adam from Kingma and Ba (2015), to maximize the conditional probability of the labels in the training dataset under the model.
@@
 SpaRTA proposes to randomly select a small subset of the model parameters to optimize while freezing the rest.
-The model parameters $\theta\in\mathbb{R}^{n}$ are partitioned into trainable $\theta_{\text{T}}\in\mathbb{R}^{m}$ and frozen $\theta_{\text{F}}\in\mathbb{R}^{n-m}$ ones, with $m\ll n$ being the number of parameters selected.
+The model parameters $\theta\in\mathbb{R}^{n}$ are partitioned into trainable $\theta_{\text{T}}\in\mathbb{R}^{m}$ and frozen $\theta_{\text{F}}\in\mathbb{R}^{n-m}$ ones, with $m\ll n$ being the number of selected parameters.
 This allows our approach to have a drastically lower memory footprint than FT by reducing the size of the optimizer state as well as faster training by reducing the number of gradients to compute.
-In our SpaRTA’s implementation [^1] https://github.com/IBM/SpaRTA, we introduce: (i) non-trainable indices $\phi\in\mathbb{R}^{m}$ containing the indices of the randomly selected elements in $\theta_{\text{PT}}$, and (ii) trainable parameters $\Delta_{\phi}\in\mathbb{R}^{m}$ representing the subset of $\Delta$ that SpaRTA learns at indices $\phi$.
+This is similar with respect to LoRA in that both reduce the memory footprint by optimizing a small set of parameters (the adapter) while freezing most (SpaRTA) or all (LoRA) of the model parameters.
+In our SpaRTA implementation [^1] Code is available at https://github.com/IBM/SpaRTA., we introduce: (i) non-trainable indices $\phi\in\mathbb{R}^{m}$ containing the indices of the randomly selected elements in $\theta_{\text{PT}}$, and (ii) trainable parameters $\Delta_{\phi}\in\mathbb{R}^{m}$ representing the subset of $\Delta$ that SpaRTA learns at indices $\phi$.
 The pseudocode for SpaRTA is given in Algorithm 1.
@@
 Finally, inference is performed, similarly to LoRA, by merging the fine-tuned $\Delta$ into $\theta_{\text{PT}}$ and then making the necessary forward passes.
-Thus, SpaRTA does not introduce any additional inference latency when deploying the resulting fine-tuned model.
+Thus, SpaRTA does not introduce any additional inference latency.
 5 Memory Usage
-Recall that SpaRTA freezes $n-m$ ($m\ll n$) of the model parameters.
+Recall that SpaRTA freezes nmn-m ($m\ll n$) of the model parameters.
 We define sparsity as $s=1-m/n\in(0,1)$, the percentage of frozen model parameters (e.g., if $1\%$ of parameters are trainable, then the sparsity is $99\%$).
-Subsequently, density is defined as $k=m/n=1-s$, the percentage of trainable model parameters.
+Subsequently, density is defined as k=m/n=1sk=m/n=1-s, the percentage of trainable model parameters.
 In practice, for a chosen sparsity $s$, one can freeze a model parameter with probability $s$, expecting a total sparsity percentage of $s$ over all model parameters.
-Thus, in expectation, $k=m/n$ percent of the model parameters are chosen as trainable, for a total of $n\,k=m$ trainable parameters.
-For SpaRTA, only $\Delta_{\phi}$ (of size $m$) is trainable, which is significantly smaller than the total number of trainable parameters for FT since $m\ll n$.
-However, the indices of these randomly chosen model parameters must be recorded into $\phi$, adding to the memory requirements.
+Thus, in expectation, $k=m/n$ percent of the model parameters are chosen as trainable, for a total of nk=mn\,k=m trainable parameters.
+For SpaRTA, only $\Delta_{\phi}$ (of size $m$) is trainable, which is significantly smaller than the total number of model parameters since $m\ll n$.
+However, the indices of these randomly chosen parameters must be recorded into $\phi$, adding to the memory requirements.
 Indices can be stored in $16$-bit integers for all the Transformer models (Vaswani et al., 2017) considered in this paper.
@@
 Thus, SpaRTA requires up to $m\,(2\times 16+16)=3m\times 16$ bits of extra memory to specify the index $\phi$ and delta $\Delta_{\phi}$ tensors.
-That is $3\,k$ times more memory than the original model, which requires just $n\times 16$ bits for storing its parameters.
+That is 3k3\,k times more memory than the original model, which requires just $n\times 16$ bits for storing its parameters.
 For instance, using SpaRTA on a model with sparsity 80%, 90%, 95% and 99% would require up to 60%, 30%, 15% and 3% more memory, respectively.
@@
 SpaRTA’s memory savings during training.
-Optimizing the full set of model parameters (FT) using Adam (Kingma and Ba, 2015) requires memory for the parameters, their gradients with their (adaptive) first and second moments estimates.
+Optimizing the full set of model parameters (FT) using Adam (Kingma and Ba, 2015) requires memory for the parameters, their gradients, and their (adaptive) first and second moments estimates.
 This requires $4\,n\times 16$ bits of memory when using a bfloat16 representation.
-In contrast, SpaRTA only optimizes $\Delta_{\phi}$, requiring a total of $m\,(4\times 16+2\times 16)+n\times 16$ bits of memory: (i) $m\,(4\times 16)$ bits for the statistics needed by Adam to optimize $\Delta_{\phi}$ of size $m$, (ii) $m\,(2\times 16)$ bits for $\phi$ with the indices identifying the model parameters associated with $\Delta_{\phi}$, and (iii) the PT model parameters ($n\times 16$ bits).
+In contrast, SpaRTA only optimizes $\Delta_{\phi}$, requiring a total of $m\,(4\times 16+2\times 16)+n\times 16$ bits of memory: (i) $m\,(4\times 16)$ bits needed by Adam to optimize $\Delta_{\phi}$ of size $m$, (ii) $m\,(2\times 16)$ bits for $\phi$ with the indices identifying the model parameters associated with $\Delta_{\phi}$, and (iii) the PT model parameters ($n\times 16$ bits).
 Memory savings appear then if and only if
-$m\,(4\!\times\!16+2\!\times\!16)+n\times 16$ $<4n\!\times\!16,$ (1)
-$kn\,(6\!\times\!16)$ $<3n\!\times\!16,$
-$k$ $<\frac{1}{2},$
-that is, SpaRTA is a real PEFT method iff $k<0.5$, i.e., a sparsity higher than 50% is required ($s\!>\!0.5$), making less than 50% of the model parameters trainable.
+$m\,(4\!\times\!16+2\!\times\!16)+n\times 16$ $<4n\!\times\!16$ (1)
+$kn\,(6\!\times\!16)$ $<3n\!\times\!16$
+$k$ $<0.5,$
+that is, SpaRTA is a real PEFT method iff $k<0.5$, i.e., a sparsity higher than 50% is required (s>0.5s\!>\!0.5), making less than 50% of the model parameters trainable.
 For instance, using SpaRTA on a model with sparsity $s=$ 80%, 90%, 95%, and 99% requires 45%, 60%, 67.5% and 73.5% less memory than full parameter FT, respectively, see Table 1.
@@
 Model storage Full FT SpaRTA $n$ (on disk) 0% 50% 80% 90% 95% 99% 2B 4 16 16 8.8 6.4 5.2 4.2 7B 14 56 56 30.8 22.4 18.2 14.8
-[…]
+Table 1: SpaRTA memory usage efficiency during training for $n=\text{2B}$ and 7B parameter models.
+Memory in Gigabytes (GB) for storing the parameters on disk (storage) as well as for FT with Adam on (i) the full set of parameters (Full FT) equivalent to 0% sparsity (ii) a sparse random subset (SpaRTA) for different sparsity percentages from 50% to 99%.
 6 Experimental Setup
-We next detail our experimental framework. Motivation is first given for the tasks followed by a description of models to be used and how these models are adapted (with manipulations) to the tasks.
+We next detail our experimental framework. Motivation is first given for the tasks followed by a description of models to be used and how they are adapted (with manipulations) to the tasks.
 6.1 Tasks
@@
 They are of particular interest for our experiments as they will show results for models with two different numbers of parameters.
-All models are readily available text-to-text, decoder-only transformer models with open weights, downloadable from Hugging Face [^4] https://huggingface.co/.
+All models are readily available text-to-text, decoder-only transformer models with open weights, downloadable from Hugging Face.
 Note that when using an instruction-following (it) model, inputs are formatted according to the conventions established when training the model on instructions.
@@
 W, MLP, norm 0.140 95.4%
-Table 3: Full table of loss and accuracy on the SST-2 task after applying our SpaRTA approach to different types of parameters in the Gemma 2B model, given that the same number of trainable parameters are selected.
-Results are averaged across 10 random seeds
+Table 3: Loss and accuracy on the SST-2 task after applying our SpaRTA approach to different types of parameters in the gemma-2b model, given that the same number of trainable parameters are selected.
+Results are averaged across 10 random seeds.
 gemma-2b gemma-2b-it mistral-7b mistral-7b-it method density (%) loss acc.
@@
 loss acc.
-PT (zero-shot) 0.425 92.0% 0.435 86.1% Full FT 100% 0.092 96.9% 0.107 96.3% 0.080 97.4% 0.080 97.3% SpaRTA 5% 0.096 96.7% 0.105 96.3% 0.084 97.2% 0.081 97.5% SpaRTA 0.5% 0.096 96.7% 0.103 96.4% 0.087 96.9% 0.087 97.0% SpaRTA 0.05% 0.106 96.3% 0.114 96.2% 0.080 97.3% 0.076 97.4% LoRA $\approx$0.050.101 96.5% 0.113 96.2% 0.086 97.1% 0.081 97.1%
+PT (zero-shot) 0.425 92.0% 0.435 86.1% Full FT 100% 0.092 96.9% 0.107 96.3% 0.080 97.4% 0.080 97.3% SpaRTA 5% 0.096 96.7% 0.105 96.3% 0.084 97.2% 0.081 97.5% SpaRTA 0.5% 0.096 96.7% 0.103 96.4% 0.087 96.9% 0.087 97.0% SpaRTA 0.05% 0.106 96.3% 0.114 96.2% 0.080 97.3% 0.076 97.4% LoRA $\approx$0.05% 0.101 96.5% 0.113 96.2% 0.086 97.1% 0.081 97.1%
 Table 4: Test loss and accuracy of models adapted to the IMDB dataset with different fine-tuning methods.
 We also report zero-shot performance of instruction following PT models.
-For training details see Table 16.
+For training details, see Table 19.
 Model: gemma-2b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc.
@@
 avg.
-PT (zero-shot) 1.34 40.5 1.77 43.7 0.75 60.7 1.41 48.3 2.91 30.9 0.91 64.6 0.85 0.7 0.88 68.5 44.7 SpaRTA (5%) ALL 0.17 93.7 0.54 81.0 0.14 95.0 0.26 89.0 0.33 87.2 0.35 85.3 0.39 56.2 0.36 84.4 84.0 SpaRTA (0.5%) 0.16 94.0 0.43 80.7 0.15 94.7 0.24 89.7 0.32 87.8 0.34 86.5 0.45 54.9 0.37 84.7 84.1 SpaRTA (0.037%) 0.18 93.1 0.48 76.9 0.16 94.5 0.26 88.8 0.35 86.4 0.36 84.5 0.45 54.8 0.40 83.6 82.8 LoRA (0.037%) Q,V 0.18 93.4 0.44 78.7 0.14 95.1 0.26 89.1 0.33 87.8 0.33 85.4 0.41 55.0 0.36 84.0 83.6 DoRA (0.037%) 0.18 93.4 0.45 78.0 0.15 95.6 0.26 89.0 0.33 87.8 0.33 85.5 0.41 57.1 0.36 84.4 83.8 SpaRTA (0.037%) 0.19 92.8 0.46 77.6 0.16 94.8 0.26 88.6 0.36 86.3 0.36 84.2 0.45 54.5 0.39 83.4 82.8 SpaRTA (0.037%) O,V 0.17 93.5 0.44 80.7 0.15 94.9 0.26 88.9 0.34 87.2 0.35 85.7 0.42 55.9 0.37 85.0 84.0
-Table 5: Test loss and accuracy of model gemma-2b-it adapted to GLUE and BoolQ datasets with different fine-tuning methods. For training details see Appendix E.2.
+PT (zero-shot) 1.34 59.5 1.77 56.3 0.75 60.7 1.41 48.3 2.91 30.9 0.91 64.6 0.85 0.7 0.88 68.5 44.7 SpaRTA (5%) ALL 0.17 93.7 0.54 81.0 0.14 95.0 0.26 89.0 0.33 87.2 0.35 85.3 0.39 56.2 0.36 84.4 84.0 SpaRTA (0.5%) 0.16 94.0 0.43 80.7 0.15 94.7 0.24 89.7 0.32 87.8 0.34 86.5 0.45 54.9 0.37 84.7 84.1 SpaRTA (0.037%) 0.18 93.1 0.48 76.9 0.16 94.5 0.26 88.8 0.35 86.4 0.36 84.5 0.45 54.8 0.40 83.6 82.8 LoRA (0.037%) Q,V 0.18 93.4 0.44 78.7 0.14 95.1 0.26 89.1 0.33 87.8 0.33 85.4 0.41 55.0 0.36 84.0 83.6 DoRA (0.037%) 0.18 93.4 0.45 78.0 0.15 95.6 0.26 89.0 0.33 87.8 0.33 85.5 0.41 57.1 0.36 84.4 83.8 SpaRTA (0.037%) 0.19 92.8 0.46 77.6 0.16 94.8 0.26 88.6 0.36 86.3 0.36 84.2 0.45 54.5 0.39 83.4 82.8 SpaRTA (0.037%) O,V 0.17 93.5 0.44 80.7 0.15 94.9 0.26 88.9 0.34 87.2 0.35 85.7 0.42 55.9 0.37 85.0 84.0
+Table 5: Test loss and accuracy of gemma-2b-it adapted to GLUE and BoolQ datasets with different fine-tuning methods.
+Standard errors can be found in Table 15, see Appendix D.
+For training details see Appendix E.2.
 Table 4 shows the results for IMDB, where each model is asked to classify a review as positive or negative.
-Each model is fine-tuned using the following adaptation methods: (i) Full parameter fine-tuning (Full FT) where all model parameters are optimized; (ii) SpaRTA for different density levels 5%, 0.5%, 0.05%, with the last allowing SpaRTA to have approximately the same number of trainable parameters as LoRA; (iii) LoRA with rank $r=8$ equivalent to about $0.05\%$ of trainable parameters compared to the model full parameter size.
+Each model is fine-tuned using the following adaptation methods: (i) Full parameter fine-tuning (Full FT) where all model parameters are optimized; (ii) SpaRTA for different density levels 5%, 0.5%, 0.05%, with the last allowing SpaRTA to have approximately the same number of trainable parameters as LoRA; (iii) LoRA with rank $r=8$, equivalent to about $0.05\%$ of trainable parameters compared to the model full parameter size.
 In Table 4, results for adaptation methods (rows) are sorted in order of descending density as to show the impact of decreasing the number of trainable parameters on the overall performance.
@@
 So far, SpaRTA has chosen trainable parameters from any layer of the transformer with equal probability. In contrast, LoRA concentrates adaption on the key and value self-attention weight matrices.
-Now, we investigate whether it is better to concentrate the selection of sparse, trainable parameters on a few parameters or equally across all.
-Specifically, given a budget for the number of parameters that can be trained, which parameters should we target for sparsification (while freezing the remaining) to achieve best task performance when adapting the sparsified parameters within the targeted types?
+Now, we investigate whether it is better to concentrate the selection of sparse, trainable parameters on a few model layers or equally across all.
+Specifically, given a budget for the number of parameters that can be trained, which type of parameters should we target for sparsification (while freezing the remaining) to achieve best task performance when adapting the sparsified parameters within the targeted types?
 We conduct an ablation study to answer this question.
-The Gemma 2B model is adapted to the SST-2 task using our SpaRTA approach targeting different combinations of parameters.
-We set a budget of 1.25M trainable parameters within the Gemma 2B model so that sparsity $s=99.95\%$.
+The gemma-2b model is adapted to the SST-2 task using our SpaRTA approach targeting different combinations of parameter types.
+We set a budget of 1.25M trainable parameters within the gemma-2b model so that sparsity is $s=99.95\%$.
 Results of these experiments are shown in Table 3.
@@
 In contrast, distributing our sparsification across different combinations of weight types may lead to significantly lower performance.
-Note that concentrating our sparsification over targeted parameters, although proven beneficial in terms of single-task adaption performance, can decrease performance when merging since it increases the chance of parameter interference.
+Corresponding results confirming, as done in Hu et al.
+(2022), that targeting Wo and Wv is suboptimal for LoRA can be found in Appendix D.
+Note that concentrating our sparsification over targeted parameter tensors, although proven beneficial in terms of single-task adaption performance, can decrease performance when merging since it increases the chance of parameter interference.
 7.3 GLUE and BoolQ
-Model: Mistral-7B-Instruct-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc.
+Model: mistral-7b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc.
 loss acc.
@@
 avg.
-PT (zero-shot) 2.38 23.4 2.54 33.6 0.66 66.3 0.85 57.5 2.42 35.5 0.70 70.0 0.72 25.2 0.39 85.2 49.6 SpaRTA (5%) ALL 0.11 96.0 0.26 89.8 0.12 95.9 0.22 90.7 0.23 91.6 0.32 87.3 0.36 62.6 0.24 90.8 88.1 SpaRTA (0.5%) 0.11 95.8 0.30 90.5 0.12 96.0 0.23 90.4 0.24 91.2 0.32 88.0 0.35 67.9 0.24 90.5 88.8 SpaRTA (0.048%) 0.12 95.7 0.31 89.2 0.12 96.1 0.23 90.5 0.24 91.1 0.31 88.2 0.34 67.3 0.26 89.5 88.4 LoRA (0.048%) Q,V 0.12 95.6 0.24 91.2 0.12 96.0 0.23 90.2 0.24 91.0 0.29 89.3 0.32 69.9 0.24 90.5 89.2 DoRA (0.048%) 0.13 95.0 0.27 90.9 0.11 96.4 0.23 90.2 0.25 90.9 0.28 88.7 0.32 68.2 0.24 90.5 88.8 SpaRTA (0.048%) 0.12 95.6 0.31 88.3 0.13 95.6 0.23 90.1 0.25 90.9 0.30 88.6 0.35 65.2 0.26 89.7 88.0 SpaRTA (0.048%) O,V 0.11 95.9 0.31 88.4 0.14 95.8 0.23 90.4 0.24 91.2 0.30 88.4 0.34 64.9 0.26 89.7 88.1
-Table 6: Test loss and accuracy of the mistral-7b-it model adapted to GLUE and BoolQ datasets with different fine-tuning methods. For training details see Appendix E.2.
+PT (zero-shot) 1.38 76.6 1.54 66.4 0.66 66.3 0.85 57.5 2.42 35.5 0.70 70.0 0.72 25.2 0.39 85.2 49.6 SpaRTA (5%) ALL 0.11 96.0 0.26 89.8 0.12 95.9 0.22 90.7 0.23 91.6 0.32 87.3 0.36 62.6 0.24 90.8 88.1 SpaRTA (0.5%) 0.11 95.8 0.30 90.5 0.12 96.0 0.23 90.4 0.24 91.2 0.32 88.0 0.35 67.9 0.24 90.5 88.8 SpaRTA (0.048%) 0.12 95.7 0.31 89.2 0.12 96.1 0.23 90.5 0.24 91.1 0.31 88.2 0.34 67.3 0.26 89.5 88.4 LoRA (0.048%) Q,V 0.12 95.6 0.24 91.2 0.12 96.0 0.23 90.2 0.24 91.0 0.29 89.3 0.32 69.9 0.24 90.5 89.2 DoRA (0.048%) 0.13 95.0 0.27 90.9 0.11 96.4 0.23 90.2 0.25 90.9 0.28 88.7 0.32 68.2 0.24 90.5 88.8 SpaRTA (0.048%) 0.12 95.6 0.31 88.3 0.13 95.6 0.23 90.1 0.25 90.9 0.30 88.6 0.35 65.2 0.26 89.7 88.0 SpaRTA (0.048%) O,V 0.11 95.9 0.31 88.4 0.14 95.8 0.23 90.4 0.24 91.2 0.30 88.4 0.34 64.9 0.26 89.7 88.1
+Table 6: Test loss and accuracy of the mistral-7b-it model adapted to GLUE and BoolQ datasets with different fine-tuning methods.
+Standard errors can be found in Table 16, see Appendix D.
+For training details see Appendix E.2.
+method targets gemma-2b-it mistral-7b-it gemma-2b mistral-7b LoRA Q, V 83.6 89.2 81.7 87.3 SpaRTA Q, V 82.8 88.0 78.3 86.7 SpaRTA O, V 84.0 88.1 83.2 85.6
+Table 7: Average accuracy (%) per adapted model.
+Averages are taken across 8 tasks (7 GLUE and BoolQ) based on Tables 5, 6, 13, and 14.
+SpaRTA is shown to be competitive with LoRA, particularly when targeting Wo (O) and Wv (V) self-attention matrices, which were previously shown to be optimal for SpaRTA.
+method targets gemma-2b-it mistral-7b-it gemma-2b mistral-7b LoRA Q, V 37.58 62.5 37.5 50 SpaRTA Q, V 0 0 0 31.3 SpaRTA O, V 62.5 37.5 62.5 18.8
+Table 8: Win rates (%) per adapted model, computed across 8 tasks (7 GLUE tasks and BoolQ) based on Tables 5, 6, 13, and 14.
+SpaRTA is shown to be competitive with LoRA, particularly when targeting Wo (O) and Wv (V) self-attention matrices, which were previously shown to be optimal for SpaRTA.
 MMLU gemma-2b gemma-2b-it mistral-7b mistral-7b-it loss acc.
@@
 We also report zero-shot performance of instruction following PT models.
-For training details see Table 17 in the Appendix.
-We now focus on comparing SpaRTA and LoRA methods using the 7 NLU task in GLUE tasks as well as BoolQ.
+For training details see Table 20 in Appendix E.
+We now focus on comparing SpaRTA and LoRA methods using the 7 NLU tasks in the GLUE benchmark as well as BoolQ.
 Table 5 presents the results for adapting gemma-2b-it.
-An equivalent set of results for the gemma-2b base model is given in Table 10, see Appendix D.
+An equivalent set of results for the gemma-2b model is given in Table 13, see Appendix D.
 Adaptation results (rows) are grouped by which type of parameters are targeted, with SpaRTA further ordered in descending order of density.
@@
 A similar set of results on GLUE and BoolQ is provided for mistral-7b-it in Table 6.
-Again, results for the mistral-7b base model can be found in Table 11 in Appendix D.
-Again, SpaRTA exhibits comparable performance to LoRA and DoRA for a low density of 0.0048%, where all methods have comparable numbers of trainable parameters.
+Again, results for the mistral-7b model can be found in Table 14 in Appendix D.
+SpaRTA exhibits comparable performance to LoRA and DoRA for a low density of 0.0048%, where all methods have comparable numbers of trainable parameters.
 And once again, with SpaRTA targeting Wv and Wo typically outperforming the targeting of Wq and Wv.
+To further illustrate the competitiveness of SpaRTA with LoRA, Table 7 summarizes accuracy results across the 8 datasets while Table 8 offers another perspective by considering win-rates among the competing methods for each model.
+Note that LoRA exhibits best performance overall when targeting the Wq and Wv self-attention matrices as mentioned in the original LoRA paper Hu et al.
+(2022), and hence we target these parameters when using LoRA.
+While we have already explored optimal targets for SpaRTA, we include results targeting Wq and Wv for completeness.
+Across both tables, SpaRTA (Wo, Wv) outperforms SpaRTA (Wq, Wv) in six out of the eight scenarios.
+LoRA and SpaRTA are equally performant on average across the two tables, demonstrating the claim that SpaRTA is competitive with LoRA.
 7.4 MMLU
 We have chosen MMLU because it is known for being a very challenging NLU task.
-Table 7 compares SpaRTA against LoRA on adapting each of our models to the MMLU multiple choice question answering task, where each model must predict the correct answer to a set of questions.
+Table 9 compares SpaRTA against LoRA on adapting each of our models to the MMLU multiple-choice question answering task, where each model must predict the correct answer to a set of questions.
 Solving this task requires a high level of world knowledge and problem solving skills.
@@
 These results demonstrate that a simple sparsifying scheme can offer a valid adaptation technique.
-This opens the possibility for further investigating sparse adaptation as the low rank approximation of a model’s parameter changes is not the only mechanism to provide a performant adaptation method.
+This opens the possibility for further investigating sparse adaptation as the low-rank approximation of a model’s parameter changes is not the only mechanism to provide a performant adaptation method.
 This indicates that what matters is not necessarily the adapter structure used in PEFT techniques, but rather the number of trainable parameters, that is relevant to the adaptation task at hand.
@@
 We compute all delta weight matrices for two well-known instruction-following fine-tuned models: gemma-2b-it and mistral-7b-it, see Section 6.2.
-As we can see in Tables 8 and 9, the feedforward (MLP) delta $\Delta$-matrices associated with each layer of the transformer network are all full rank [^5] We compute the rank of a matrix as the number of singular values that are greater than zero, given a specified tolerance, using torch.linalg.matrix_rank.
+As we can see in Tables 10 and 11, the feedforward (MLP) delta $\Delta$-matrices associated with each layer of the transformer network are all full rank [^4] We compute the rank of a matrix as the number of singular values that are greater than zero, given a specified tolerance, using torch.linalg.matrix_rank.
 We did some sensitivity analysis to determine such tolerance.
@@
 We demonstrated our method with varying density levels.
-The total average number of trainable parameters in each case was approximately: $99$M (5% density), $10$M (0.5%) and $800$k (0.037%) for the (base and instruct) Gemma 2B models; and $349$M (5% density), $35$M (0.5%) and $3.5$M (0.048%) for the (base and instruct) Mistral 7B models.
+The total average number of trainable parameters in each case was approximately: 9999M (5% density), 1010M (0.5%) and 800800k (0.037%) for the (base and instruct) Gemma 2B models; and 349349M (5% density), 3535M (0.5%) and 3.53.5M (0.048%) for the (base and instruct) Mistral 7B models.
 For LoRA, we factorized the changes in the query and value self-attention projection weight matrices with rank $r=8$ decomposition matrices, which were optimized while keeping all other model parameters frozen.
-The number of new trainable parameters introduced by the LoRA approach in the Gemma 2B and Mistral 7B models were approximately $925$k and $3.5$M, respectively.
+The number of new trainable parameters introduced by the LoRA approach in the Gemma 2B and Mistral 7B models were approximately 925925k and 3.53.5M, respectively.
 Also, we set $\alpha=16$ to scale the LoRA adapters.
@@
 We also noticed that larger models require higher sparsity levels to eliminate overfitting given the same training data.
-On the other hand, Head adapation always resulted in significant underfitting, making this fine-tuning method always underperform.
-We also observed that SpaRTA achieves a speedup over Full FT for sparsity levels greater than 90%.
-Thus, at some sparsity level between 80% and 90%, computing gradients on a smaller amount of trainable parameters starts to offset the extra compute of adding and subtracting the deltas before and after the forward pass respectively.
-These extra operations require slicing the original parameter tensors over the randomly selected indices.
 Appendix D Additional Experiment Results
-Tables 10 and Table 11 present the complete set of results on GLUE and BoolQ for the PT base models gemma-2b and mistral-7b, respectively.
-Tables 12 and 13 present standard errors from corresponding experiments.
+Table 12 confirms targeting Wo and Wv is not the best choice for LoRA (96.2% accuracy for Wo and Wv versus 96.4% for Wq and Wv which is the LoRA default).
+These results can be compared with the corresponding SpaRTA results in Section 7.2.
+Tables 13 and Table 14 present the complete set of results on GLUE and BoolQ for the gemma-2b and mistral-7b models, respectively.
+Tables 15 and 16 present standard errors from corresponding experiments.
+LoRA targets rank Loss Accuracy
+Wq 16 0.155 94.8 %
+Wk 24 0.182 94.5%
+Wv 24 0.116 96.2%
+Wo 16 0.114 96.2%
+Wq, Wk 8 0.152 94.8%
+Wq, Wv (default) 8 0.110 96.4%
+Wq, Wo 6 0.107 96.6%
+Wk, Wv 12 0.110 96.7%
+Wk, Wo 8 0.110 96.9%
+Wv, Wo 8 0.114 96.2%
+Wq, Wv, Wo 4 0.108 96.7%
+Wk, Wv, Wo 6 0.113 96.7%
+Wq, Wk, Wv, Wo 3 0.108 96.8%
+MLP 1 0.112 96.2%
+Table 12: Test loss and accuracy on SST-2 after applying LoRA to different types of parameters in the gemma-2b model, given that the same number of trainable parameters are selected.
+Results averaged across 10 random seeds.
 Model: gemma-2b QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ Method Targets loss acc.
@@
 For training details see Appendix E.2.
-Model: Mistral-7B-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method Targets loss acc.
+Model: mistral-7b QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ Method Targets loss acc.
 loss acc.
@@
 loss mcc loss acc.
-PT (zero-shot) 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 SpaRTA (5%) ALL 0.00 0.14 0.03 1.54 0.01 0.51 0.00 0.21 0.01 0.26 0.01 0.48 0.00 0.88 0.01 0.54 SpaRTA (0.5%) 0.00 0.16 0.03 0.64 0.01 0.46 0.00 0.17 0.00 0.13 0.01 0.07 0.01 0.75 0.01 0.25 SpaRTA (0.037%) 0.00 0.12 0.01 0.55 0.00 0.17 0.00 0.18 0.00 0.14 0.00 0.45 0.01 0.44 0.00 0.09 LoRA (0.037%) Q,V 0.00 0.11 0.01 1.10 0.00 0.35 0.00 0.03 0.00 0.11 0.00 0.22 0.02 1.40 0.01 0.12 DoRA (0.037%) 0.00 0.11 0.01 0.55 0.00 0.20 0.00 0.03 0.00 0.09 0.00 0.32 0.01 1.31 0.01 0.29 SpaRTA (0.037%) 0.00 0.08 0.00 1.04 0.01 0.11 0.00 0.00 0.00 0.26 0.00 0.22 0.02 2.83 0.00 0.41 SpaRTA (0.037%) O,V 0.00 0.13 0.01 0.64 0.00 0.27 0.00 0.16 0.00 0.17 0.02 0.22 0.01 4.36 0.00 0.22
+SpaRTA (5%) ALL 0.00 0.14 0.03 1.54 0.01 0.51 0.00 0.21 0.01 0.26 0.01 0.48 0.00 0.88 0.01 0.54 SpaRTA (0.5%) 0.00 0.16 0.03 0.64 0.01 0.46 0.00 0.17 0.00 0.13 0.01 0.07 0.01 0.75 0.01 0.25 SpaRTA (0.037%) 0.00 0.12 0.01 0.55 0.00 0.17 0.00 0.18 0.00 0.14 0.00 0.45 0.01 0.44 0.00 0.09 LoRA (0.037%) Q,V 0.00 0.11 0.01 1.10 0.00 0.35 0.00 0.03 0.00 0.11 0.00 0.22 0.02 1.40 0.01 0.12 DoRA (0.037%) 0.00 0.11 0.01 0.55 0.00 0.20 0.00 0.03 0.00 0.09 0.00 0.32 0.01 1.31 0.01 0.29 SpaRTA (0.037%) 0.00 0.08 0.00 1.04 0.01 0.11 0.00 0.00 0.00 0.26 0.00 0.22 0.02 2.83 0.00 0.41 SpaRTA (0.037%) O,V 0.00 0.13 0.01 0.64 0.00 0.27 0.00 0.16 0.00 0.17 0.02 0.22 0.01 4.36 0.00 0.22
 Table 15: Standard errors of test loss and accuracy of model gemma-2b-it adapted to GLUE and BoolQ datasets with different fine-tuning methods.
-Model: Mistral-7B-Instruct-v0.3 QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc.
+Model: mistral-7b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ Method targets loss acc.
 loss acc.
@@
 loss mcc loss acc.
-PT (zero-shot) 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 SpaRTA (5%) ALL 0.00 0.09 0.01 0.24 0.01 0.26 0.00 0.03 0.00 0.08 0.01 0.47 0.01 1.91 0.00 0.16 SpaRTA (0.5%) 0.00 0.05 0.01 1.07 0.01 0.18 0.00 0.10 0.00 0.09 0.02 0.25 0.01 1.28 0.00 0.09 SpaRTA (0.048%) 0.00 0.10 0.03 0.55 0.00 0.20 0.00 0.03 0.00 0.11 0.01 0.32 0.02 1.87 0.00 0.20 LoRA (0.048%) Q,V 0.00 0.16 0.01 0.60 0.01 0.44 0.00 0.06 0.00 0.13 0.01 0.33 0.01 1.16 0.00 0.17 DoRA (0.048%) 0.02 0.76 0.02 0.87 0.00 0.17 0.00 0.03 0.00 0.07 0.00 0.17 0.01 1.22 0.00 0.16 SpaRTA (0.048%) 0.00 0.06 0.01 0.73 0.00 0.23 0.00 0.04 0.00 0.05 0.01 0.43 0.01 1.17 0.00 0.27 SpaRTA (0.048%) O,V 0.00 0.02 0.01 0.42 0.01 0.34 0.00 0.07 0.00 0.13 0.02 0.24 0.01 0.97 0.00 0.23
+SpaRTA (5%) ALL 0.00 0.09 0.01 0.24 0.01 0.26 0.00 0.03 0.00 0.08 0.01 0.47 0.01 1.91 0.00 0.16 SpaRTA (0.5%) 0.00 0.05 0.01 1.07 0.01 0.18 0.00 0.10 0.00 0.09 0.02 0.25 0.01 1.28 0.00 0.09 SpaRTA (0.048%) 0.00 0.10 0.03 0.55 0.00 0.20 0.00 0.03 0.00 0.11 0.01 0.32 0.02 1.87 0.00 0.20 LoRA (0.048%) Q,V 0.00 0.16 0.01 0.60 0.01 0.44 0.00 0.06 0.00 0.13 0.01 0.33 0.01 1.16 0.00 0.17 DoRA (0.048%) 0.02 0.76 0.02 0.87 0.00 0.17 0.00 0.03 0.00 0.07 0.00 0.17 0.01 1.22 0.00 0.16 SpaRTA (0.048%) 0.00 0.06 0.01 0.73 0.00 0.23 0.00 0.04 0.00 0.05 0.01 0.43 0.01 1.17 0.00 0.27 SpaRTA (0.048%) O,V 0.00 0.02 0.01 0.42 0.01 0.34 0.00 0.07 0.00 0.13 0.02 0.24 0.01 0.97 0.00 0.23
 Table 16: Standard errors of test loss and accuracy of model mistral-7b-it adapted to BLUE and BoolQ datasets with different fine-tuning methods.
 Appendix E Training Hyper-Parameters
-The hyper-parameters used for investigating SpaRTA and other adaptation methods on IMDB and GLUE are summarized next.
+The hyper-parameters used for investigating SpaRTA and other adaptation methods are summarized next.
 E.1 IMDB
@@
 That is 77%, 75%, 73%, and 71% of the original training data of 25,000 examples, respectively.
-Model: google/gemma-2b-it method QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ SpaRTA (5%) 1e-5 5e-4 5e-5 5e-5 1e-5 2e-4 5e-5 5e-5 SpaRTA (0.5%) 2e-4 1e-3 2e-4 2e-4 2e-4 5e-4 1e-3 2e-4 SpaRTA (0.037%) 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 LoRA (0.037%) 2e-4 1e-3 5e-4 5e-4 5e-4 2e-4 5e-4 5e-4 DoRA (0.037%) 2e-4 1e-3 5e-4 5e-4 5e-4 2e-4 5e-4 5e-4 SpaRTA (0.037%) 5e-4 1e-3 5e-4 5e-4 5e-4 5e-4 1e-3 5e-4 SpaRTA (0.037%) 5e-4 1e-3 5e-4 5e-4 5e-4 5e-4 1e-3 5e-4
+Model: gemma-2b-it Method QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ SpaRTA (5%) 1e-5 5e-4 5e-5 5e-5 1e-5 2e-4 5e-5 5e-5 SpaRTA (0.5%) 2e-4 1e-3 2e-4 2e-4 2e-4 5e-4 1e-3 2e-4 SpaRTA (0.037%) 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 1e-3 LoRA (0.037%) 2e-4 1e-3 5e-4 5e-4 5e-4 2e-4 5e-4 5e-4 DoRA (0.037%) 2e-4 1e-3 5e-4 5e-4 5e-4 2e-4 5e-4 5e-4 SpaRTA (0.037%) 5e-4 1e-3 5e-4 5e-4 5e-4 5e-4 1e-3 5e-4 SpaRTA (0.037%) 5e-4 1e-3 5e-4 5e-4 5e-4 5e-4 1e-3 5e-4
 Table 17: Optimal learning rates for gemma-2b-it model for different fine-tuning methods.
-Model: Mistral-7B-Instruct-v0.3 method QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ SpaRTA (5%) 1e-5 5e-6 5e-6 1e-5 1e-5 5e-5 5e-5 5e-6 SpaRTA (0.5%) 2e-4 1e-4 5e-5 5e-4 2e-4 5e-4 2e-4 5e-5 SpaRTA (0.048%) 5e-4 5e-4 5e-4 1e-3 5e-4 1e-3 1e-3 1e-3 LoRA (0.048%) 2e-4 1e-4 2e-4 2e-4 1e-4 2e-4 2e-4 1e-4 DoRA (0.048%) 2e-4 1e-4 2e-4 2e-4 1e-4 2e-4 2e-4 1e-4 SpaRTA (0.048%) 2e-4 1e-4 2e-4 5e-4 2e-4 1e-3 1e-3 5e-4 SpaRTA (0.048%) 2e-4 1e-4 2e-4 5e-4 2e-4 1e-3 1e-3 5e-4
+Model: mistral-7b-it Method QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ SpaRTA (5%) 1e-5 5e-6 5e-6 1e-5 1e-5 5e-5 5e-5 5e-6 SpaRTA (0.5%) 2e-4 1e-4 5e-5 5e-4 2e-4 5e-4 2e-4 5e-5 SpaRTA (0.048%) 5e-4 5e-4 5e-4 1e-3 5e-4 1e-3 1e-3 1e-3 LoRA (0.048%) 2e-4 1e-4 2e-4 2e-4 1e-4 2e-4 2e-4 1e-4 DoRA (0.048%) 2e-4 1e-4 2e-4 2e-4 1e-4 2e-4 2e-4 1e-4 SpaRTA (0.048%) 2e-4 1e-4 2e-4 5e-4 2e-4 1e-3 1e-3 5e-4 SpaRTA (0.048%) 2e-4 1e-4 2e-4 5e-4 2e-4 1e-3 1e-3 5e-4
 Table 18: Optimal learning rates for mistral-7b-it model for different fine-tuning methods.
@@
 Table 19: (IMDB) Training parameters used with each fine-tuning method and model in Table 4.
-$\ast$ in number of epochs indicates early stopping was used.
+An $\ast$ in number of epochs indicates early stopping was used.
 For SpaRTA, parameters for density 5%, 0.5% 0,05% are reported.
@@
 For GLUE and BoolQ, training was done using a simple grid search over the learning rates: [1e-3, 5e-4, 2e-4, 1e-4, 5e-5, 1e-5, 5e-6].
-The optimal learning rates found are given in Tables 14 and 15.
+The optimal learning rates found are given in Tables 17 and 18.
 Weight decay was set to 0.